A Complete Bibliography of Publications in the ACM Symposia on Theory of Computing (STOC) for

1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

05 April 2013 Version 2.07

Title word cross-reference

2 [Mil84]. 2n [BJ85]. 3 [Blu89, Yao83]. = [DL80]. ^{1} [Bar86, BT87]. d [YY85]. g [HI87]. K [Ko88, CSY84, GKS86]. l [DGS88]. n [Mey85, adH83]. O(((m + n)n^{2} + (m + n)^{1.5}n)L) [Vai87]. O(1) [FN89]. O(g) [HI87]. O(lg n) [Bra85]. O(log log N) [AKSS86]. O(log N) [Upf89, BP84]. O(n) [Bro86]. O(n^{0.4}) [Blu89]. O(n^{2} log n log^{∗} n) [Blo80]. O(n log n) [AKS83]. ω [CH81, SV89]. s [BKRU89]. t [BKRU89].

-approximation [Blu89]. -automata [SV89]. -coloring [Blu89]. -connected [Mil84]. -dimensional [Mey85, YY85, adH83]. -exclusion [DGS88]. -hulls [CSY84]. -page [GKS86]. -sequence [CH81]. -space [Yao83].

2n [Cop84].

Abstract [Ang80, BL82, Bro80, CD80, Fag80, HVY81, HS80, Hof80, Imm82, JY81, KT80, KR81, Kin81, KLLM81, Lei82, Llo80, MS82, OvL80, Rei80, Sto80, Str80, Var82, Wil82, Yao81a, Yao82, VW84]. accelerating [CV86]. Access [BMS81, Bro80, HLR87, Mye83]. achieve [AK88]. achieving [DHS84]. ACM [ACM80, ACM81]. acyclic [BFM^{+}81]. adaptive [ABNLP89b]. addition [Huy85, Kos82]. adjusting [ST83]. admit [Tom80b]. affine [JM85]. agreement [DS82, FM88]. agreements [GY89]. algebra [BOKR84, FR85]. Algebraic [Smo87, BO83, Can88, Cle88, EM80, Hua84, Ja’80, KLL84, vzG83]. algebraically [Ier89]. algebraically-closed [Ier89]. algebras [PP81, Pra80, Ron87]. algorithm [AM84, AGSS87, AA87, And85, BBL83, BOFKT86, BO88, Blo80, Blu89, BFF85, BLS87b, Cla85, CL86, DF89, FM80, FK84, GT83, Har85, Kar84, KW84, LV86, Lub85b, May81, MR87, Mul86, Orl88, SS86, TV86, Vai87, Wig82, adH83]. algorithmic [Mir80]. Algorithms [LM85, AC88, Awe87, Awe89, Bac82, Bac87, BS85, CEGS89, Cla84, Cla87, CM89, CV86, DS87, DS82, FG88, FM88, FL89, GT88a, GW88, Gal84, Kan83b, KV86, KR88, KT80, LS86, MMS88, Mey85, Mot89, NNS87, PKR82, PY88b, Pnu83, RS81, Tom80b, Tur83, vzG83]. allocation [CL86, Lyn80]. Almost [GT88a, GK86, Hås86, JWR81]. Almost-optimum [GT88a]. ALOGTIME [Bus87]. alternating [ITT82, Kin81, MT82]. Alternation [Kan83a]. Amortized [ST84]. Analysis [HLR87, Bac87, GM81, HMT88b, KLMS84, LS86, LM88, Mot89, PY88b, Tur83]. Angeles [ACM80]. Annual [ACM81, ACM80, ACM82, ACM83, ACM84, ACM85, ACM86, ACM87, ACM88, ACM89]. answering [Yao82]. any [BM88, GMW87]. application [CHMP81, Kal87, SvEB84]. applications [BFM88, CEGS89, Cla86, GW88, Har84, LS86, NY89, Yao83]. approach [Cyp80, Fed89, Fre87, GT86, PSS80, Sip83b, YY85, óSY83]. approximate [FG88, LV86, Sto83a, Wig82]. approximated [PW89]. approximating [BYE82, DF89]. Approximation [Cla87, BBL83, Blu89, Bro86, Cla84, GT87, HS84b, JS88, PY88a]. approximations [Raz89]. April [ACM80, ACM83, ACM84]. arbitrary [Mul86]. archimedean [Vit84]. architecture [PY88b]. architecture-independent [PY88b]. area [BP84, Blu84]. area-maximum [Blu84]. arithmetic [Alt84, BK80, Bus85, CU89, CW87, JY81, Kos86, Vai87]. arrangement [EG86]. array [AK82, GR85]. arrays [KA86, RT85]. Aspects [Sie86, SS89]. assemble [BL82]. associative [CFL83a]. asynchronous [AFL81, HV88, Iwa83]. atomic [DGS88]. Automata [VW84, Eng83, Gác83, GH82, KV89, MS81b, RS89b, SV89, Ukk80]. automaton [DG81]. auxiliary [Pip80]. Average [CM84, AM84, BDCG89, LS86, Lev84, Mot89]. average-case [Mot89]. axiomatization [Mir80, SU80].

backoff [HLR87]. backtrack [CSW85]. Bandwidth [MS81a, Tur83]. based [HMT88b, OSS84]. be [CS80, GK86, PW89]. behavior [BJL^{+}84, BS85]. Berkeley [ACM86]. better [DG81, DG82, MSS87, MS82]. between [AM84, AEL85, AFL81, Bop89, CHMP81, DKS^{+}81, KL80b, KA86, LTT89, PU88, Upf84]. bidirected [Gab83]. biinfinis [NP82]. bin [BJL^{+}84, KLMS84]. binary [BK80, Cul85, HVY81, JWR81, PRS81, ST83]. bipartite [GT88a]. bits [BOCS83, KLL84]. bivariate [Kal82]. blocking [FFP86]. blue [JWK81]. bodies [DF89]. boolean [Alt84, BP87, Bus87, CK89, KLPV87, KV89, Nat87, Smo87]. Borel [Sip83a]. Boston [ACM83]. bottleneck [HS84b]. bound [KZ88]. Bounded [Bar86, DS89, AM84, BGM82, Bop84, Bus85, DPPU86, FSS83, HS84a, Mil80, Mir80, RST82, Sim81]. Bounded-width [Bar86]. Bounds [BDFP83, CD82, PRS81, ABH^{+}86, BH87, BO83, BRT89, DPGR83, DGS84, Fic81, Fic83, FadHRW85, Hås86, Kin88, Lei84, LS86, LY86, LS81, Maa84, PKR82, PSS80, Smo87, Val83, VS85]. branch [KZ88]. branch-and-bound [KZ88]. branching [ABH^{+}86, Bar86, BDFP83, EH82, ES84]. breaking [Adl83, GPS87]. broadcast [Gre82]. Building [BBE^{+}84]. Butterfly [BCHR88]. Byzantine [Bra85, FM88, GY89].

calculus [KM81]. California [ACM80, ACM82, ACM86]. can [CS80, GK86]. canceling [GT88b]. cannot [PW89]. Canonical [BL83]. capture [Imm83]. captures [CRRS89]. cascades [CV86]. case [BDCG89, CM84, DGS88, GT83, GS88, LS86, Mot89]. cell [FS89b]. cellular [Gác83, GR85]. certain [Kal87]. certified [GK86]. chains [JS88]. Channel [Mir84]. channels [HLR87]. characterization [BMS81, Old83]. characterize [Ven87]. check [BK89]. Chervonenkis [BEHW86]. Chicago [ACM88]. chip [BK80]. chordal [NNS87]. circuit [Alt84, Sip83a, Smo87]. Circuits [Yao89, ACR88, AUY83, AIS84, Bop84, CFL83a, Fic83, Hås86, KW88, Kis82, MT87, Rei89, Sie86, Val83]. circulations [GT88b]. City [ACM87]. class [BMS81, SU80]. Classes [HVY81, Eng83, Imm83, KL80b, PY88a, Sim81]. Classifying [BEHW86]. clauses [Fag80]. clock [DHS84]. closed [Ier89, SS89, Sim81]. closure [Kal86, Tom80b]. clustering [FG88]. coin [Cle86, CV86]. coins [GS86]. collapses [Hem87]. collections [FK84]. coloring [Blu89, VL88, Wig82]. combinatorial [AIK81, Yan88]. combinators [CEGS89]. common [Kal85]. communicating [Vaz85]. Communication [OE84, PS82, ABNLP89a, DGS84, FL84, Fur87, Gre82, HMT88a, HR88, Iwa83, LTT89, RS81, VB81, Vaz85]. commutative [Huy85]. commute [CRRS89]. Compact [ABNLP89b]. compactness [KA84]. Comparative [Pip80]. Comparison [Alt84, KS84]. comparisons [BJ85]. Competitive [MMS88]. complement [Sim81]. complementation [Rei82]. Complete [Mir80, MS81a, BBE^{+}84, Lev84, PRS81, SU80, USS81]. Completeness [BOW88]. Complexity [EM80, AIK81, ABNLP89a, BDCG89, BOKR84, BP87, BK80, CSW85, hC81, CM81, CEG87, DGS84, Eng83, FT85, For87, FS89b, Fur87, GMR85, Gre82, HMT88a, HV86, Huy85, IL80, Imm83, JWK81, Jw80, Kal87, KL80b, Kin88, Kre86, Lei84, Lei81, MT87, Orl81, PS82, PY82, PY88a, Sim81, Sip83a, Sip83b, SC82, Smo87, Sto83a, Var82, Vaz85]. compose [BKP84]. composition [LR86]. compositional [Ger84]. Compression [GS85]. computable [BMS81, Imm82, Kal86]. computation [BO83, BOW88, BC80, BH82, CD80, CM81, FadHRW85, FS89a, Gác83, GS83, HL89, JY81, Kin81, LY86, PSS80, Upf84, Vis84, Yao81b, Yao89]. computational [Cla86, FR85, MT87, RS89a]. computationally [KUW85a]. computations [BP87, Can88, RST82, Yao81a, ABO84]. compute [CD82, HS80, Mul86]. computer [GP81, JY80]. computers [BS88, SS86]. Computing [ACM80, ACM81, ACM82, ACM83, ACM84, ACM85, ACM86, ACM87, ACM88, ACM89, Cle88, FB86, Kal85, AGSS87, HVY81, MS82]. concepts [BEHW86]. concrete [Bac82]. concurrency [dBMOZ85a, Yan81, dBZ82, dBMOZ85b]. Concurrent [Pel85, Bea86, Bop89, CVW86, Man84]. concurrent-write [Bea86, Bop89]. concurrrent [DS89]. Conductance [JS88]. cone [Hof80]. Conference [ACM80, ACM81]. configurations [OvL80]. conflicts [Wil84]. conjecture [Bir89, Bsh89, KMR89]. conjectures [LPS86]. connected [AK82, KA86, Mil84, SS86]. connection [KLMR89]. connections [KL80b]. connectivity [BKRU89, KW88, KT80]. connectors [DDPW83]. consequences [All87, IR89]. considerations [Vaz87]. consistency [DL80]. consistent [PW89]. constant [ABO84, Cle88]. constrained [FW88, Gab83, MS81a]. constraints [CK85, Llo80, OE84]. constructible [DS89, Ist88]. Constructing [KUW85b, PU87, Sei86, FK84]. constructive [CU89]. consumption [ACR88, Kis82]. contraction [AH82]. control [Yan81]. controlled [LLS87]. convergence [DM84]. converse [Str81]. Convex [Cha81, AGSS87, CEG87, DEY86, DF89, KV86, Sei86]. convexity [KR81]. Coordinate [GPS89]. core [GL89, SvEB84]. Correcting [DMMU84]. correctness [Yan81]. cost [GT87, GT88b, Orl88, Sei86, Sto80]. counter [hC81, DG81]. counting [Awe87, Sto83a]. cover [BYE82, CRRS89]. CRCW [BH87]. CREW [Nis89]. Critical [Llo80]. Cryptographic [DLM82, KV89, BOCS83, BOW88, HS85, LR86, NY89]. cryptography [Kil88]. cryptosystems [Adl83]. cutting [CEG87]. cycle [Fic81, Mil84]. cycles [BFF85, CM89, GT88b, Ist88, IS87, KS88b].

Data [Fre83, FS84, ABNLP89b, DSST86, FS89b, ST81]. database [BFM^{+}81, CK85, CGKV88, Fag80, Yan81]. databases [SU80]. DC [ACM84]. Deadlock [Tou80]. Deadlock- [Tou80]. Decidability [Kos82]. Decidable [CGKV88, CH81, Fei83, KL80a]. Deciding [ES84]. Decision [EH82, BH87, FL89, KUW85a, KV87, MT82, Nis89, Ukk80]. decomposability [Iwa83]. decompositions [Cha81]. Definability [MP80, dR84]. degree [DPPU86, Gab83]. degree-constrained [Gab83]. delay [Agg85]. deletion [AH82]. Denotational [dBZ82]. dense [Wil82]. dependencies [CLM81, Fag80, KCV83, SU80]. depth [AA87, AAK89, ABO84, Alo85, Alt84, Bop84, DDPW83, Hås86, KW88, KPPY84]. depth-first [AAK89]. Designing [BK89, CV86]. desirable [Upf84]. Detailed [OvL80]. Detecting [KS88b, CM89, VV85a]. Detection [CD80, Fic81]. Determining [Gon84, BOFKT86, FM80]. determinism [MS82]. Deterministic [AKSS86, AKS87, CV86, BO88, DT83, Für82, GS88, Jw80, Maa84, Ukk80, Upf89]. DFA [PW89]. diagram [AGSS87]. diagrams [GS83]. diameter [DF83]. difference [AFL81]. different [HR88]. difficult [CS80, FB86, Ron87]. Digital [KR81, KA84]. dimension [AM84, BEHW86]. dimensional [GR85, ITT82, Mey85, Sei86, YY85, adH83]. direct [Bsh89]. directed [AAK89]. discreet [LW83]. discrete [LLS87, LW83]. disjoint [BS85, GT83, PU87]. disks [KA84]. distance [STT86]. Distributed [RS81, Vit84, Awe87, ABNLP89b, Awe89, BOW88, CDDS85, Lyn80, MS82, PKR82]. distribution [Pla80]. division [Kos86, Rei89]. division-free [Kos86]. divisors [BMS84, Kal85]. DOL [CH81]. domain [BL85]. dominators [Har85]. double [LM88]. Doubly [Lub85a]. Draft [GM81, Vit82]. Dual [AEL85]. Dynamic [MP80, MT87, Pra80, CL86, CM89, Fei83, FH82, FS89b, KS88b, Man84, Orl81, Pel85, ST81, Wil82, Yao80, Str81]. Dynamically [OvL80].

easier [Cha84, CD80]. easy [HS80, MVV87, VV85a]. Edge [AH82, Blu84, PRS81]. edge-contraction [AH82]. Edge-deletion [AH82]. edgings [Koz84]. effect [Cul85]. Efficiency [Vaz87, AFL81, FL89, PU88, ST84]. Efficient [BFDS84, PR85, Yao80, Awe84, CL86, Gab83, GP81, KU86, LV86, OSS84, SvEB84]. eigenvalue [BGM82, FKS89]. Eighteenth [ACM86]. electing [FL84]. election [Awe87]. elections [Vit84]. electrical [CRRS89]. elementary [BOKR84]. elements [Bea89]. elimination [Ier89]. ellipsoids [Pos84]. embeddable [Lic80]. Embedded [CLM81]. embedding [Sto80]. embeddings [dFPP88]. emulations [KLMR89]. encryption [GM82]. ends [MS81b]. Energy [ACR88, Kis82]. Ensembles [NP82]. entropic [Yao81a]. environment [Man84, Wil82]. Equational [CK85, Koz84]. Equations [CHMP81, BL85, Kos86, OSS84]. equivalence [CH81, Gon84, Huy85, IL80, Ukk80]. equivalent [KUW85a]. era [DGS88]. errors [KL88]. estimate [Bac82]. Ethernet [GGMM85, Wil84]. Euler [AIS84]. Evaluating [Cop84]. evaluation [Kos86]. Every [KS84]. exact [GMS84]. exactly [Bar86, Ko88]. Examples [KM81]. exchange [Blu83, KLLM81]. exclusion [DGS88]. execution [Str80]. existence [All87]. expander [PU87]. Expanders [Alo85, JM85, LPS86]. Expanding [Mot89]. expected [BJL^{+}84, Blo80, BS85, Bra85, Cla84, KT80]. expected-time [Cla84, KT80]. Explicit [LPS86]. Exponential [Val83, GPS89, Hem87]. exponentiating [FT85]. Expressing [Yan88]. expressions [Gon84]. Expressiveness [AC89, EH82]. Exptime [HSI83]. Extended [Ang80, BL82, Bro80, CD80, Fag80, HVY81, HS80, Hof80, Imm82, KT80, KR81, Kin81, KLLM81, Lei82, Llo80, MS82, Rei80, Sto80, Str80, Var82, VW84, Wil82, Yao81a, Yao82, Bsh89, Bac82]. extensions [PP81]. extremal [BDDG82]. extremely [Pnu83]. face [Sei86]. facets [PY82]. factor [Kal87]. Factoring [Len83, BMS84, Val89]. Factorization [Hua84, Bac83, Kal82, KLL84]. fails [KMR89]. fair [CVW86, Hen84, Pnu83]. faithful [PP81]. familiar [Tom80b]. fan [CFL83a]. fan-in [CFL83a]. fary [dFPP88]. Fast [Bac82, Cla84, HL89, JY81, KV86, Lyn80, Mey85, NNS87, BOFKT86, KW84, Mul86, PSS88, Val89]. faster [Orl88]. Fault [DPPU86, Fel85, BOW88, BS88, BFDS84, DHSS84]. fault-tolerant [BOW88, BS88]. faults [DMMU84, HLN87]. faulty [Cle86, HL89]. feasible [Upf84]. feasibly [CU89]. field [Ier89, Mul86]. fields [AL86, FT85, Hua84, Hua85, Len83]. fifteenth [ACM83]. file [BBE^{+}84]. files [BBE^{+}84, Wil82]. Finding [AL86, AIS84, BJ85, BDDG82, GT88b, Mil84, Bea89, BFF85, GMS84, Har85, Hua85, PKR82]. fine [BT87]. fingers [BFG85]. Finite [BT87, AL86, FT85, HS84a, Hua84, Hua85, KV89, Len83, RS89b, dR84]. firing [CDDS85]. first [ACM89, AA87, AAK89, Yap80]. fixed [Fed89, FM80, KLMR89]. fixed-connection [KLMR89]. flips [Cle86]. flow [Gab83, GT86, GT87, Har85, Orl88, Sie86, Str80]. flows [KV86, SNS85]. Fooling [DG81]. Forests [GW88]. formal [Iwa83]. forms [Che81, FSS83]. formula [Bus87]. formulae [KLPV87, KV89, KPPY84, Pla80]. formulas [Cle88]. foundation [Kis82]. Founding [Kil88]. Four [Yan86]. fourteenth [ACM82]. fragments [Bus85, DL80]. frames [GW88]. Francisco [ACM82]. free [CS80, Kos86, Tou80]. function [BM88]. Functional [CU89]. functions [AM84, Alt84, BMS81, Bop84, CFL83a, CD82, GL89, HVY81, ILL89, Kal86, KLD86, Lev85, NY89, Nat87, SvEB84]. Further [Cla86].

game [GMW87, JWK81]. games [AIK81, GW88, GH82]. general [AAK89, Bea89, BC80, GP81, GS83, Har84, May81, YY85]. Generalized [FJ80, Adl83, DDPW83]. generalizers [DDPW83]. generals [Bra85]. generate [Bac83]. generating [FK84, Vaz85]. generation [ILL89]. generators [All87, Lev85, LR86]. genus [FM80, HI87, Mil80]. geometric [BEHW86, Can88, Cla84, YY85]. Geometry [Vai88, BOKR84, Cla86, GBT84, RS89a, STT86]. GF [Cop84]. given [BM88, Kal85]. global [Ang80, KS89]. good [KS84, Lei82, RT85]. Graph [AK82, CS80, CRRS89, HMT88a, Kin88, KLLM81, VL88, Wig82]. Graphs [JWR81, AAK89, BGM82, BL83, BYE82, BFF85, BS84, CM89, FM80, Fre87, FKS89, FSS83, GKS86, GPS87, Har85, HI87, HS84b, Hof80, IS87, KNR88, KS88b, Lic80, Mil80, Mil84, Mot89, MS81b, NNS87, PU87, RT85, SS87, Sto80, Str80, SNS85, Yan86, dFPP88]. graphy [MR87]. greatest [Kal85]. grid [LS86, Sto80]. group [Bab85]. groups [BLS87a, DF83].

half [Cle86]. Hamilton [BFF85]. hard [BN89, Bro86, GL89, KM81, Mey85, VL88]. hard-core [GL89]. hash [NY89, SvEB84]. hashing [FNSS88, KU86, LM88, SS89]. having [Ko88]. help [Wil85]. helps [Vai88]. Hensel [Kal87]. Heuristics [SPR80]. hiding [AFK87]. hierarchical [AACS87]. hierarchies [Ko88, RST82]. hierarchy [Bus85, Cai86, Für82, Hem87, Mye83]. higher [KV87, Sei86]. higher-dimensional [Sei86]. higher-order [KV87]. Highly [BGSV89]. Hilbert [CHS85]. Hoare [Old83]. homing [RS89b]. honest [RBO89]. Horn [Fag80]. hulls [CSY84, Sei86]. hyperbolic [STT86]. hypercube [HLN87]. hypercubes [HL89]. hypothesis [Hua85, Kur82, Bac82].

I/O [JWK81]. identity [FFS87]. Illinois [ACM88]. impact [FL84]. Imperfect [LLS87]. implementation [KU86]. implementations [Bro80, EM80, Tom80b]. implicational [CLM81]. Implicit [FN89, KNR88]. impossibility [DHS84]. improbability [GY89]. Improved [Kan83b, VS85]. inclusion [KCV83]. Independence [JY80, FK84]. independent [KW84, Lub85b, PY88b, Pla80]. induction [Pra80]. inequalities [CHS85, Yao80]. inequality [LMR86]. Inference [RS89b, CLM81, KCV83]. infinitary [dBMOZ85a, dBMOZ85b]. infinite [Har84, IS87, PSS88]. infinity [FadHRW85]. information [AFK87, AUY83, GM82, Man84, PSS80, Sie86]. information-theoretic [PSS80]. initial [JY81]. instance [Lev84]. instances [VL88]. integer [AEL85, Kan83b, Rei89, Val89]. integers [Bac83]. integral [Kal82]. interactive [BOGKW88, BFM88, GMR85, GS86]. internal [FV85]. interpolation [BO88]. interpretations [CU89, Mir80]. interprocess [RS81]. Intersecting [Cha84]. intersection [FS84]. intervals [FK84]. intractability [BOGKW88]. Introducing [LV86]. invariance [SvEB84]. inversion [MVV87]. inverted [BBE^{+}84]. Investigations [BS88]. irreducible [AL86]. Island [ACM85]. ismorphism [KMR89]. Isomorphism [BGM82, Lic80, Mil80, FM80, Hof80]. isomorphisms [BL85, KLD86]. Issues [Yan81]. Iterated [Eng83]. iterative [PP81].

Kannan [Bir89]. keeping [GM82]. key [Adl83, Mun87]. keys [Blu83]. knapsack [Adl83, GMS84, Yao81b, adH83]. knowledge [BFM88, FFS87, For87, GMR85, HV86, HMT88b, HV88]. knowledge-based [HMT88b]. known [Bac83]. Kraft [Bro80]. labeling [BL83]. LALR [USS81]. Languages [Imm83, dBMOZ85a, Bar86, BL85, Orl81, Var82, dBMOZ85b]. large [SU80]. lattice [Kan83b]. law [CK89]. layout [Blu84, FP80, Lei82]. layouts [KLLM81, LM85]. leader [Awe87, FL84]. learnability [KLPV87]. learnable [BEHW86, Val84]. Learning [KL88, KV89, Nat87]. length [Blu84, BRT89, PRS81]. level [AIK81, EM80]. levels [Ko88]. lexical [Lub85a]. lifting [Kal87]. like [Old83]. likelihood [HR83]. limitations [KV89, Yao81a]. limited [Alo85, DDPW83]. Limits [Bea86, Cle86, IR89]. line [Fre83, FS84, IL80, Kal85, Kal87, MSS87, MMS88, PSS80]. Linear [KT80, LMR86, AGSS87, AEL85, BBE^{+}84, FS89a, GT83, GM81, GS88, Har85, Kar84, PR85, RV83, SC82, TV86, Vai87, Yan88, adH83]. linear-time [GT83, TV86]. Linearity [LN88]. linearly [HS85]. Lines [CEGS89]. linked [Die82]. list [Bro80, Die82, DS87, ST84]. lists [Kos81]. livelock [Tou80]. livelock-free [Tou80]. Liveness [DM84]. Local [Ang80, KS89, Yao89]. Localized [Kos81]. locking [Yan81]. Log [Wil84, LW83, LMR86]. Log-logarithmic [Wil84]. logarithmic [AIS84, KW88, RV83, Sei86, Wil84]. logarithms [Cop84]. LOGCFL [Ven87]. Logic [MP80, BKP84, CHMP81, CGKV88, EH82, ES84, FV85, Fei83, FH82, Ger84, HR83, Koz84, MS81b, Old83, Pel85, SV89, Str81]. Logics [Rei80, HS84a, SC82, VW84, VS85]. logspace [BN89, AKS87]. logspace-hard [BN89]. look [DHSS84]. looping [Str81]. Low [AIK81]. Lower [BO83, BRT89, DGS84, Fic81, Kin88, LS81, ABH^{+}86, DPGR83, FadHRW85, Hås86, LY86, Maa84, PKR82, Smo87, Val83, VS85]. machine [Sim81]. machines [Bea86, BL82, Bop89, hC81, DG81, DG82, DT83, GKS86, ITT82, Maa84, MSS87, Vit82, BMS81]. macro [CV86]. Maintaining [Die82, Wil82, DS87, Man84, OvL80]. majority [RBO89]. Making [DSST86]. malicious [KL88]. Manhattan [BBL83]. manipulation [GS83]. many [CS80]. Markov [JS88]. marriages [Fed89]. marry [Bro86]. Massachusetts [ACM83]. Matching [MVV87, GT88a, GS81, Gal84, KUW85b, LV86, LS86, Sto83b, SPR80, Vai88]. matchings [Mot89]. matrices [Lub85a]. Matrix [CW87, Mul86, MVV87]. matroid [FS84, GW88]. matroids [RS80]. maximal [And85, KW84, Lub85b]. maximum [Blu84, GT86, PKR82]. maximum-finding [PKR82]. May [ACM81, ACM82, ACM84, ACM85, ACM86, ACM87, ACM88, ACM89, BKP84, GS88]. measure [KA84, Sto80]. Measures [Kin81]. Measuring [Kis82]. median [BJ85]. memory [AACS87, AC88, DMMU84, KU86, RS82]. mental [GMW87, GM82]. merging [BH82, HS82]. mesh [AK82, KA86, SS86]. mesh-connected [AK82, KA86]. method [Raz89, Ukk80]. methods [Smo87]. metric [DM84]. metrical [BLS87b]. micro [CV86]. Milwaukee [ACM81]. minimal [Fel85]. minimax [LS86, PRS81]. minimization [Tur83]. Minimum [Pos84, Awe87, BP84, Cla84, FK84, Fre83, GT87, GT88b, Orl88, PW89]. minimum-cost [GT87, GT88b]. mixing [JS88]. modal [FV85, VW84, VS85]. model [AACS87, BC80, CM81, Yap80]. Modelling [Hen84]. models [Agg85, BH82, BL85, HS84a, JY81, Upf84]. modes [HR88]. monoids [BT87]. Monotone [KW88, Bop84, KPPY84, Val83]. motion [Cla87, FW88, óSY83]. motion-planning [óSY83]. mots [NP82].

Multi [BOGKW88, CFL83b]. Multi-party [CFL83b]. Multi-prover [BOGKW88]. Multicommodity [SNS85, KV86]. multicounters [Vit82]. multidimensional [KLMS84]. multikey [FNS^{+}88]. Multiparty [BN89, CD88, RBO89]. multiple [DS82, HLR87]. multiplexed [Iwa83]. multiplication [CW87]. multiplicity [BGM82]. multivariate [BO88, Kal82, Len83].

n2 [Spi83]. nature [Pra80]. NC [AA87, BLS87a, Bar86, BT87, CM89, KUW85b, VV85b]. nearby [Lyn80]. necessary [Yan86]. negative [GT88b]. net [May81]. network [AKS83, Awe84, BP84, DHSS84, Fel85, Gab83]. networks [Ang80, BS88, BP87, BFDS84, DPPU86, Fed89, FFP86, HS82, KLMR89, RV83, Tou80, BCHR88]. nexptime [HSI83]. nineteenth [ACM87]. no [Tom80b]. node [Sto80]. Non [BFM88, FFP86, FNSS88, BOW88, DGS88, Ukk80]. non-atomic [DGS88]. Non-blocking [FFP86]. non-cryptographic [BOW88]. Non-interactive [BFM88]. Non-oblivious [FNSS88]. non-real-time [Ukk80]. nondeterminism [Kan83a]. Nondeterministic [GS88, DG82, GKS86, Maa84, MT82, Mir80]. nonlinear [DPGR83, GS88]. nonrandomness [KLL84]. nontrivial [GKS86]. nonuniform [KL80b]. Normal [FSS83, Che81]. note [KLD86]. Notes [HS82]. notions [AUY83]. NP [DL80, HSI83, MS81a, VV85a, You83]. NP-Complete [MS81a]. NP-P [HSI83]. number [AM84, Cle88, DL80, HI87, Hua84, Pla80]. numbers [BMS84, KLL84, KR88].

O [JWK81]. oblivious [FNSS88, Gol87, Kil88, KPU88, Vit82]. obtained [JM85]. off [MSS87]. off-line [MSS87]. office [Cla85]. on-line [Fre83, FS84, MMS88, PSS80]. once [DMMU84, RS82]. One [FadHRW85, Lev85, Cai86, CK89, DG81, DG82, GKS86, GMS84, GL89, IR89, ILL89, KLD86, Maa84, MSS87, NY89, Vit82]. one-tape [GKS86, Maa84, Vit82]. One-way [Lev85, GL89, IR89, ILL89, KLD86, NY89]. online [BLS87b]. operations [Vai87]. optima [AEL85]. Optimal [Awe87, BH87, BCHR88, Bop89, DKS^{+}81, FG88, FM88, FP80, Gal84, HMM85, KA86, Rei89, BLS87b, GS81, Hås86, SS86, Tom80a]. Optimization [PY88a, CGKV88, Kre86, Orl81, Yan88]. optimum [GT88a]. oracle [AFK87, Cai86, Kur82, KMR89]. orbit [KL80a]. order [Die82, DS87, GPS89, KS89, KV87, MS81b, Yap80]. orderings [Lub85a]. orders [KMK89]. organizing [CHS85]. orientation [Spi83]. orthogonal [Vai85].

P [DL80, HSI83, Kal86]. P-computable [Kal86]. packet [Tou80, Upf89]. packing [BJL^{+}84, KLMS84]. page [GKS86, HI87]. pagenumber [BS84]. pages [Yan86]. pairs [Fre87]. papers [ACM80, ACM81]. Parallel [AAK89, GPS87, KU86, Kos86, vzG83, AKSS86, And85, AIS84, Bea86, BOFKT86, Bop89, BH82, CV86, CD82, DT83, Fic83, FadHRW85, FT85, GP81, Gal84, KW84, KZ88, Lei84, LY86, Lub85b, MT87, Mul86, NNS87, PR85, PY88b, Upf84, VB81, Vis84, Yao81b]. parallelism [AK88, Kin81, LV86]. parallelizable [BGSV89]. parallelization [MR87]. parameter [Lei81]. partial [GM82, KMK89]. partition [Yao83]. party [CFL83b]. Pascal [Old83]. Pascal-like [Old83]. passing [Lei81]. path [And85, Blo80, Cla87, Fel85, Llo80]. paths [Awe89, Cyp80, Fre87, PU87, SS84, SNS85]. PDL [Koz83]. pebble [JWK81]. pebbles [CS80, KS88a, Wil85]. pebbling [CS80, Pip80]. Pebblings [Koz84]. perfect [BMS84, For87, KUW85b, SvEB84, SPR80]. performance [SS89]. periodic [IS87]. permanent [Bro86, JS88]. Permutation [BLS87a, DF83, LR86]. permutations [IR89]. persistent [DSST86]. Petri [May81]. planar [BYE82, BS84, Fre87, KS89, LM85, Mil84, Sto80, SNS85, Yan86, dFPP88]. plane [Lic80, OvL80]. Planning [FW88, Cla87, óSY83]. play [GMW87, GM82]. point [Fed89]. poker [GM82]. Polling [RS89a]. poly [LMR86]. poly-log [LMR86]. polygon [AGSS87]. polygons [BDDG82, TV86]. polyhedra [Cha81]. polyhedral [SS84]. polymorphic [Lei81]. Polynomial [DS82, FR85, Ist88, KLL84, AH87, Bar86, BOFKT86, BO88, BMS81, Bus85, Cai86, CM89, DF89, FL89, FM80, Gon84, Imm82, JY81, Kal82, KCV83, Kar84, KLD86, Ko88, KS88b, LM83, Orl88, PW89, Tom80b, You83, adH83]. polynomial-size [Bar86]. polynomial-time [Cai86, CM89, FL89, FM80, Kar84, KLD86]. polynomials [AL86, FT85, HS80, Hua84, Kal85, Kal87, Len83]. polytopes [DEY86]. polytypes [CEG87]. poset [KS84]. possibility [DHS84]. post [Cla85]. power [Bea86, Fur87, KS88a, Kan83a]. powerful [HS84b]. Powers [HS84b]. PRAM [BH87]. PRAMS [Nis89]. precision [HMM85]. predicate [GL89]. prefix [BP87, Fic83]. prehension [BFG85]. Preliminary [AK82, DG81, FP80, FJ80, GS81, GM81, Gre82, HS82, JWR81, JY80, JY81, Kos82, MT82, PKR82, Pip80, Pip82, RS82, Sim81, Vit82, PSS80, Tom80a]. presence [HLN87, HMM85, KL88]. presented [ACM80, ACM81]. preserving [KLMR89]. primality [PSS88]. primes [AH87, GK86, Hua84, PSS88]. Primitives [GS83]. privacy [CK89]. Private [GS86]. Probabilistic [FS89a, GM82, HS84a, MT82, Pip82, Tur83, ABO84, Ber86, Cla85, Fei83, FH82, KLMS84, Koz83, Pnu83, Rei80, RST82, Sim81, Upf84]. probabilities [KV87]. probability [Cai86]. probe [FN89, FS89b]. Probing [DEY86, GM81]. problem [And85, Bus87, CLM81, Cla85, CDDS85, CH81, Cyp80, Fic81, FL84, GMS84, GT86, Huy85, IL80, KL80a, KW84, KV87, LN88, Lub85b, May81, PW89, Tom80a, VV85b, VL88, Yao81b, adH83]. Problems [Lev84, AH82, AK82, Awe87, BH87, BGSV89, CSY84, CGKV88, DL80, FR85, Gab83, GBT84, GT88a, GT87, Har85, HS84b, Ja’80, Jw80, KCV83, Kan83b, KT80, KLMS84, Kre86, MMS88, Mey85, MS81a, Mot89, MS81b, Orl81, Yan88, Yap80, vzG83]. procedure [KZ88]. procedures [EH82, Lei81, Old83]. Proceedings [ACM82, ACM83, ACM84, ACM85, ACM86, ACM87, ACM88, ACM89, ACM80, ACM81]. process [CHMP81]. processes [Hen84, LLS87]. processor [AK82, DS82, Llo80, VV85b]. processors [Ang80, Cle86, KA86, Vit84]. program [Mir80]. programming [Kan83b, KV86, Kar84, LMR86, Rei80, Vai87, Yao80]. programs [AEL85, ABH^{+}86, Bar86, BK89, BDFP83, CGKV88, CVW86, IL80, JY81, Kal85, KUW85a, Old83, VW84, VS85, Yan88, Yap80]. progressions [CW87]. projective [Lic80]. Proof [Bir89, GMR85, GS86]. proof-systems [GMR85]. proofs [BOGKW88, FFS87]. Properties [BFM^{+}81, Ven87, Ang80, DM84, Ger84, HMT88a, Kal86, Kin88, KV87, Mir80, You83]. property [JS88]. Propositional [Str81, Fei83, KM81, SC82]. protection [Gol87]. protocol [Awe84, Bra85]. protocols [BN89, CFL83b, CD88, DLM82, HLR87, RBO89, Wil84]. Provable [BL85, IR89]. Provably [RT85, Val89, CL86, Lei82]. prover [BOGKW88]. Providence [ACM85]. proving [PKR82]. Pseudo [ILL89, LR86]. Pseudo-random [ILL89, LR86]. pseudorandom [All87, BN89, KR88, Lev85]. PSPACE [Cai86, Can88, USS81]. PSPACE-complete [USS81]. public [Adl83, GS86]. purpose [GP81]. Pushdown [MS81b, DG81, Eng83, Ukk80]. pushdowns [Pip80].

quadrangle [Yao80]. Quadratic [Maa84, AM84, KV86, OSS84, Val89]. Quantifier [Ier89]. quasi [Val89, Vaz85]. quasi-random [Vaz85]. quasi-uniform [Val89]. queries [AC89, Imm82, Vai85, Yao82, YY85]. query [Var82]. quickly [GK86].

R [VV85b]. R-NC [VV85b]. R. [Bir89]. radicals [LM83]. radio [ABNLP89a]. RAM [CD82]. Ramanujan [LPS86]. RAMs [Gol87]. Random [RS80, VL88, AH87, AA87, Bac83, BFF85, Bro86, Cai86, Cla86, DF89, FKS89, Gon84, ILL89, KUW85b, Kur82, KMR89, LLS87, LR86, Mye83, SS87, Vaz85, Vaz87, BMS81]. Randomized [Awe89, KR88, Vis84, Bac87, Bra85, KZ88, RS89a]. randomness [Bab85, Fur87, KPU88, Sip83b]. range [Vai85, Yao82]. rank [Mul86]. ranking [FJ80, GS85]. rapid [JS88]. reachability [Kos82, May81, MS81b]. reaching [GY89]. Real [Vit82, BOFKT86, GR85, RS81, Ukk80]. Real-time [Vit82, GR85]. Realistic [Bac87]. reason [Ger84, HR83]. Reasoning [CVW86, HV88, HV86]. recognize [Bar86]. Recognizing [AH87]. Reconfiguring [HLN87]. reconnaissables [NP82]. rectangles [DKS^{+}81, FK84, GY80]. rectilinear [Ber86]. recursive [AC89]. red [JWK81]. red-blue [JWK81]. reducibilities [You83]. reduction [Gab83, Kal82]. regions [FK84]. registers [Cle88]. regular [CHMP81, FKS89, PP81, RT85]. related [Awe87, CM81, CSY84, DL80, GBT84, GT88a, HS85, Har85, Kan83b, Mot89]. relation [Upf84]. Relational [Imm82, SU80, Var82]. relationship [AEL85]. relative [KMR89]. Relativized [Ko88]. Reliable [Gác83]. remove [BOGKW88]. reorientation [KS89]. repeated [Che81]. Report [GS81]. representation [GPS89, KNR88]. require [GS88, KW88]. requires [BJ85, GPS89, Vai87]. residues [Val89]. resistance [CRRS89]. resolved [JS88]. resolving [Wil84]. resource [Llo80]. resources [Lyn80]. respect [Alt84]. restricted [AC89, KPPY84, Val83]. restrictions [Mey85]. result [Har84]. results [BJL^{+}84, Ber86, CM81, JY80]. Retraction [óSY83]. reuse [RS82]. Reversal [hC81]. rewriting [Che81]. Rhode [ACM85]. Riemann [Bac82, Hua85]. ring [FL84, Vit84]. roots [BOFKT86, Hua85]. Rotation [STT86]. rounds [Alo85, Bra85]. routability [LM85]. Routing [BH82, ABNLP89b, BBL83, DHSS84, KPU88, LM85, Mir84, PU88, RT85, Tom80a, Upf89]. routings [BFDS84, Fel85]. RSA [BOCS83]. rules [ST84].

sampling [Cla86, RS89a]. San [ACM82]. Scaling [GBT84]. scheduling [Llo80, VV85b]. schematology [Pip80]. scheme [OSS84, Upf89]. schemes [BFM^{+}81, GMY83, Gre82, Mir80, VB81]. Science [JY80]. search [AA87, AAK89, CHS85, Cul85, FL89, FN89, KUW85a, Kos81, adH83]. searches [CSW85]. Searching [Mun87, FNS^{+}88]. Seattle [ACM89]. second [FKS89, MS81b]. second-order [MS81b]. secrecy [OE84]. secret [Blu83, GM82, RBO89]. secure [CD88]. security [BOCS83, Cle86, HS85]. segments [JY81]. selection [AKSS86, CM84, FJ80]. Self [CHS85, ST83]. Self-adjusting [ST83]. Self-organizing [CHS85]. semantics [dBMOZ85a, FV85, dBZ82, dBMOZ85b]. semaphore [Wil84]. semi [Vaz87]. semi-random [Vaz87]. semigroups [Huy85]. separates [Cai86]. separations [Bop89]. separators [GKS86, Mil84]. sequence [CH81]. sequences [BN89, BRT89, Ist88, RS89b, Vaz85]. sequential [Bea89, BC80, CHS85, Wil82]. serial [LV86]. set [BBE^{+}84, BS85, GT83, GY80, KW84, LN88, Lub85b]. sets [FK84, HSI83, PSS88, Sip83a, You83, dFPP88]. seventeenth [ACM85]. shared [KU86]. sharing [RBO89]. shortest [Awe89, Blo80, Cla87, Fre87, SS84, SNS85]. shortest-path [Blo80]. shuffle [KLLM81]. shuffle-exchange [KLLM81]. shuffled [Iwa83]. sign [BM88]. signature [Car82, GMY83, OSS84]. Simple [Ron87, CD82, GR85, Lub85b, Mil84, TV86]. simplex [AM84]. simulation [AKS87, Gol87, Vit82]. simulations [BCHR88, GKS86, KA86, Pip82]. Single [Kal87, BOCS83, Mun87]. Single-factor [Kal87]. sixteenth [ACM84]. Size [BP87, Bar86, RV83, Rei89]. Size-time [BP87]. slightly [Vaz85]. slightly-random [Vaz85]. Small [dFPP88, Hås86, Mil84, Val89]. smaller [AM84]. smoothness [AK88]. software [Gol87]. solution [GMS84, PR85, Tom80a]. solutions [FS84, FR85, VV85a]. Solvability [LM83]. Solving [GT87]. Some [All87, BJL^{+}84, Can88, KL80b, You83, Bac87, FR85, Ja’80, Jw80, KLL84, PY82, Ukk80]. sort [GM81, RV83]. sorted [Kos81]. sorting [AKS83, Alo85, BP84, BC80, BH82, Cha84, Lei84, SS86]. sources [LLS87, Vaz85, Vaz87]. Space [RST82, Sim81, Vai85, Yao82, Yap80, Bea89, BC80, BKRU89, CEGS89, GS81, GS88, Ja’80, Jw80, LTT89, PU88, SvEB84, Tom80b, Yao83]. Space-bounded [RST82, Sim81]. space-combinators [CEGS89]. Space-time [Vai85, Yao82, Yap80]. spaces [DM84, SS84]. spanning [Awe87, Cla84, Fre83, Pos84]. Sparse [HSI83, BO88, GPS87]. special [GT83]. specifications [BKP84, EM80]. spectra [SS87]. speed [GT88a, Vis84]. speed-ups [GT88a, Vis84]. Speedups [DT83]. squad [CDDS85]. stability [GGMM85]. Stable [BFG85, Fed89]. stamp [DS89]. Steiner [Ber86]. steps [AM84]. storage [Bro80, CL86, GPS89]. store [DG81]. Storing [FNS^{+}88]. straight [IL80, Kal85, Kal87]. straight-line [IL80, Kal85, Kal87]. straightness [KR81]. strategy [Lei82]. string [GS81, Gal84, LV86]. strings [Iwa83]. Strong [GMY83, Hem87, Vaz85]. Strongly [CM89, Orl88]. structural [You83]. structure [BT87, IS87, ST81]. structures [ABNLP89b, DSST86, Fre83, FS84, FS89b, dR84]. subdivisions [GS83]. subgraph [Gab83]. sublinear [GS88, Tom80b]. subpolynomial [Agg85]. substantially [GS88]. successive [GT87]. successor [dR84]. sufficient [Yan86]. sum [Bsh89]. Sums [BMS84, GW88]. super [KW88]. super-logarithmic [KW88]. Superconcentrators [DDPW83, Alo85]. supporting [dFPP88]. sweeping [EG86]. switching [Tou80]. Symmetric [Rei82, Huy85]. symmetry [GPS87]. symmetry-breaking [GPS87]. Symposium [ACM80, ACM81, ACM82, ACM83, ACM84, ACM85, ACM86, ACM87, ACM88, ACM89]. synchronization [Awe84, DHS84]. synchronizing [RS81]. synchronous [AFL81, DT83, FL84]. system [Lyn80]. systems [AFL81, dBMOZ85a, BLS87b, Che81, CH81, DS89, GMR85, GS86, HV88, Huy85, Kos82, Llo80, PR85, dBMOZ85b]. table [FNS^{+}88, Mun87]. tables [PU88]. tape [GKS86, Maa84, SvEB84, Vit82]. tapes [DG82, MSS87]. task [BLS87b, Llo80]. tasks [GS88]. tautologies [KM81]. technique [Gab83, HS84b, PKR82, RS89a]. techniques [CV86, GBT84, VW84]. temporal [BKP84, EH82, Ger84, HS84a, SV89, SC82]. term [Che81]. terms [CHMP81]. test [DGS88]. Testing [HS80, Hof80, IS87, Car82, LM85, Mil80, USS81]. tests [PSS88]. text [BBE^{+}84]. their [AEL85, BK89, CLM81, NY89]. theorem [ABO84]. theorems [BOW88]. theoretic [PSS80, Sip83b, VW84]. theories [CK85]. Theory [ACM80, ACM81, ACM82, ACM83, ACM84, ACM85, ACM86, ACM87, ACM88, ACM89, Bab85, BDCG89, Car82, DL80, Gol87, Ier89, Pla80, Smo87, Val84, Vaz85]. Thirteenth [ACM81]. those [Bar86]. three [BFG85, FadHRW85, GR85]. three-dimensional [GR85]. Threshold [Bop84, SS87]. Tight [Lei84, LS86, Für82]. Time [GS81, Ja’80, AH87, AGSS87, AKSS86, AIS84, Bea89, BMS81, BP84, BP87, Blo80, BBE^{+}84, BC80, BKRU89, Cai86, Cla84, CM89, CD82, DS89, DF89, EH82, ES84, FL89, FM80, FS89a, FR85, Für82, GT83, GR85, Gon84, Gre82, GS88, HV86, HV88, Imm82, Iwa83, JY81, KCV83, Kar84, KT80, KLD86, Ko88, KS88b, KPU88, LM83, LMR86, PSS80, RS81, RV83, Spi83, TV86, Tom80b, Ukk80, Vai85, Vit82, Yao82, Yap80]. time-multiplexed [Iwa83]. time-randomness [KPU88]. Time-space [Ja’80, Bea89, BC80]. Time-space-optimal [GS81]. time-stamp [DS89]. times [CRRS89]. tolerance [DPPU86, Fel85]. tolerant [BOW88, BS88, BFDS84, DHSS84]. Topological [Sto83b]. Topologically [EG86]. topology [KS89]. tossing [CV86]. tradeoff [Bea89, Blu84, BC80, KPU88, PU88, Yao82].

Tradeoffs

[Agg85, LTT89, Ja’80, Vai85, Yap80].

Trading [Bab85, BKRU89].

transcendental [KLL84]. transfer

[AUY83, Kil88]. transformations [JM85]. Transition [dBMOZ85a, Ger84, dBMOZ85b]. Transitive [Spi83, Tom80b]. translating [GY80]. trapdoor [BM88]. traversal [BRT89]. traversing [Ist88]. treatment [Iwa83, Pnu83]. tree [Awe87, BL82, FP80]. Trees [GH82, BO83, Ber86, Cla84, Cul85, Fre83, Har84, HVY81, JWR81, Kin81, MT82, Nis89, PRS81, ST81, ST83]. triangulating [TV86]. triangulations [STT86]. triconnectivity [MR87]. trivalent [FSS83]. truncated [HS85]. Turing [GKS86, ITT82, Maa84, MSS87, Sim81, Vit82]. twelfth [ACM80]. twentieth [ACM88]. twenty [ACM89]. twenty-first [ACM89]. Two [ABH^{+}86, Ber86, DS87, DG82, DPGR83, ITT82, MSS87, PSS88, Tom80b, AM84, BDFP83, DG81, FadHRW85, LMR86, Mun87, VV85b, Vaz85]. Two-dimensional [ITT82]. two-processor [VV85b]. two-way [DG81]. typed [BL85]. types [GPS89].

Unary [KCV83]. Unbounded [CFL83a]. uncertainty [HMM85]. unconditionally [CD88]. undirected [BKRU89, SNS85]. unexpected [BJL^{+}84]. Uniform [Kal86, dR84, dBMOZ85a, KL80b, Val89, dBMOZ85b]. union [BS85, GT83, LN88]. Unique [Che81, Iwa83, Bea89, VV85a]. Universal [NY89, VB81, BRT89, Ist88]. university [SS89]. unprovability [LN88]. update [ST84]. updates [Cul85]. updating [Fre83, FS84]. upper [VS85]. ups [GT88a, Vis84]. Using [AK88, HL89, RS89b, Vaz87, Yao80].

valence [FSS83]. value [Bus87]. Vapnik [BEHW86]. variables [Che81, HS85, LMR86]. Vector [Str80, Huy85, Kos82]. Vegas [MS82]. Verifiable [RBO89]. Verifying [KMK89]. Version [AK82, DG81, FP80, FJ80, Gre82, HS82, JWR81, JY80, Kos82, MT82, PKR82, Pip80, Pip82, RS82, Sim81, PSS80, Tom80a]. versus [GS86, HSI83, SvEB84]. vertex [BYE82]. via [CW87]. Virtual [AC88]. VLSI [Agg85, ACR88, AUY83, BP84, Blu84, Car82, CM81, Kis82, Lei82, LM85, LS81, MS82, Mir84, Sie86, Yao81a]. volume [DF89, FB86]. Voronoi [AGSS87, GS83].

Washington [ACM84, ACM89]. way [DG81, Ger84, GL89, IR89, ILL89, KLD86, Lev85, NY89]. weak [JY81]. weight [Awe87]. weighted [SPR80]. which [HS80, Imm83, Lei82, Tom80b, Vai87]. White [Wil85, KS88a]. whose [AM84]. width [Bar86, BDFP83]. wire [Tom80a]. wire-routing [Tom80a]. wiring [DKS^{+}81]. Wisconsin [ACM81]. within [PW89, RS81]. Work [KLMR89, BK89, GS88]. Work-preserving [KLMR89]. write [Bea86, Bop89, DMMU84, RS82]. write-once [DMMU84, RS82].

York [ACM87].

Zero [FFS87, BFM88, CK89, For87, GMS84, HMT88b]. zero-knowledge [BFM88, For87]. zero-one [CK89, GMS84].

References

Aggarwal:1987:RNA

[AA87] A. Aggarwal and R. Anderson. A random NC algorithm for depth first search. In ACM [ACM87], pages 325–

334. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/

stoc/28395/p325-aggarwal/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p325-aggarwal/p325-
aggarwal.pdf. ACM order no.

508870.

Aggarwal:1987:MHM

[AACS87] A. Aggarwal, B. Alpern, A. Chandra, and M. Snir. A model for hierarchical memory. In ACM [ACM87], pages 305–

314. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p305-aggarwal/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p305-aggarwal/p305-
aggarwal.pdf. ACM order no.

508870.

Aggarwal:1989:PDF

[AAK89] A. Aggarwal, R. J. Anderson, and M.-Y. Kao. Parallel depthfirst search in general directed graphs. In ACM [ACM89], pages 297–308. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p297-aggarwal/
p297-aggarwal.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/73007/p297-

aggarwal/. ACM order no. 508890.

Ajtai:1986:TLB

[ABH^{+}86] M. Ajtai, L. Babai, P. Hajnal, J. Komlos, and P. Pudlak. Two lower bounds for branching programs. In ACM [ACM86], pages 30–38. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p30-ajtai/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/12130/p30-
ajtai/p30-ajtai.pdf. ACM

order no. 508860.

Alon:1989:CRC

[ABNLP89a] N. Alon, A. Bar-Noy, N. Linial, and D. Peleg. On the complexity of radio communication. In ACM [ACM89], pages 274–285. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p274-alon/p274-
alon.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p274-alon/. ACM

order no. 508890.

Awerbuch:1989:CDD

[ABNLP89b] B. Awerbuch, A. Bar-Noy, N. Linial, and D. Peleg. Compact distributed data structures for adaptive routing. In ACM [ACM89], pages 479–

489. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p479-awerbuch/
p479-awerbuch.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/73007/p479-

awerbuch/. ACM order no. 508890.

Ajtai:1984:TPC

[ABO84] Miklos Ajtai and Michael Ben-Or. A theorem on probabilistic constant depth Computations. In ACM [ACM84], pages 471–474. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Aggarwal:1988:VMA

[AC88] Alok Aggarwal and Ashok Chandra. Virtual memory algorithms. In ACM [ACM88], pages 173–185. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p173-aggarwal/
                                        ; http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        62212/p173-aggarwal/p173-
                                        aggarwal.pdf. ACM order no.

508880.

Afrati:1989:ERR

[AC89] F. Afrati and S. S. Cosmadakis. Expressiveness of restricted recursive queries. In ACM [ACM89], pages 113–

126. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p113-afrati/p113-
                                        afrati.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/73007/p113-afrati/. ACM

order no. 508890.

ACM:1980:CPT

[ACM80] ACM, editor. Conference proceedings of the twelfth annual ACM Symposium on Theory of Computing: papers presented at the symposium, Los Angeles, California, April 28–30, 1980.

ACM Press, New York, NY, USA, 1980. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

ACM:1981:CPT

[ACM81] ACM, editor. Conference proceedings of the Thirteenth Annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Milwaukee, Wisconsin, May 11–13, 1981. ACM Press, New York, NY, USA, 1981. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

ACM:1982:PFA

[ACM82] ACM, editor. Proceedings of the fourteenth annual ACM Symposium on Theory of Computing, San Francisco, California, May 5–7, 1982. ACM Press, New York, NY, USA, 1982. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

ACM:1983:PFA

[ACM83] ACM, editor. Proceedings of the fifteenth annual ACM Symposium on Theory of Computing, Boston, Massachusetts, April 25–27, 1983. ACM Press, New York, NY, USA, 1983. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

ACM:1984:PSA

[ACM84] ACM, editor. Proceedings of the sixteenth annual ACM Symposium on Theory of Computing, Washington, DC, April 30–May 2, 1984. ACM Press, New York, NY, USA, 1984. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

ACM:1985:PSA

[ACM85] ACM, editor. Proceedings of the seventeenth annual ACM Symposium on Theory of Computing, Providence, Rhode Island, May 6–8, 1985. ACM Press, New York, NY, USA, 1985. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13 1985. ACM order no. 508850.

ACM:1986:PEA

[ACM86] ACM, editor. Proceedings of the Eighteenth annual ACM Symposium on Theory of Computing, Berkeley, California, May 28–30, 1986. ACM Press, New York, NY, USA, 1986. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986. ACM order no. 508860.

ACM:1987:PNA

[ACM87] ACM, editor. Proceedings of the nineteenth annual ACM Symposium on Theory of Computing, New York City, May 25–27, 1987. ACM Press, New York, NY, USA, 1987. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. ACM order no. 508870.

ACM:1988:PTA

[ACM88] ACM, editor. Proceedings of the twentieth annual ACM Symposium on Theory of Computing, Chicago, Illinois, May 2– 4, 1988. ACM Press, New York, NY, USA, 1988. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988. ACM order no. 508880.

ACM:1989:PTF

[ACM89] ACM, editor. Proceedings of the twenty-first annual ACM Symposium on Theory of Computing, Seattle, Washington, May 15–17, 1989. ACM Press, New York, NY, USA, 1989. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989. ACM order no. 508890.

Aggarwal:1988:ECV

[ACR88] Alok Aggarwal, Ashok Chandra, and Prabhakar Raghavan. Energy consumption in VLSI circuits. In ACM [ACM88], pages 205–216. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p205-aggarwal/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p205-aggarwal/p205-
aggarwal.pdf. ACM order no.

508880.

aufderHeide:1983:PLS

[adH83] Friedhelm Meyer auf der Heide. A polynomial linear search algorithm for the ndimensional knapsack problem. In ACM [ACM83], pages 70– 79. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Adleman:1983:BGK

[Adl83] Leonard M. Adleman. On breaking generalized knapsack public key cryptosystems. In ACM [ACM83], pages 402–412. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Aharoni:1985:DIL

[AEL85] R. Aharoni, P. Erdős, and N. Linial. Dual integer linear programs and the relationship between their optima. In ACM [ACM85], pages 476–

483. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p476-aharoni/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p476-aharoni/p476-aharoni.

pdf. ACM order no. 508850.

Abadi:1987:HIO

[AFK87] M. Abadi, J. Feigenbaum, and J. Kilian. On hiding information from an oracle. In ACM [ACM87], pages 195–

203. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p195-abadi/;
http://www.acm.org/pubs/articles/
proceedings/stoc/28395/p195-
abadi/p195-abadi.pdf. ACM

order no. 508870.

Arjomandi:1981:DEB

[AFL81] Eshrat Arjomandi, Michael J.

Fischer, and Nancy A. Lynch. A difference in efficiency between synchronous and asynchronous systems. In ACM [ACM81], pages 128–132. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Aggarwal:1985:TVM

[Agg85] A. Aggarwal. Tradeoffs for VLSI models with subpolynomial delay. In ACM [ACM85], pages 59–68. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p59-aggarwal/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p59-aggarwal/p59-aggarwal.

pdf. ACM order no. 508850.

Aggarwal:1987:LTA

[AGSS87] A. Aggarwal, L. Guibas, J. Saxe, and P. Shor. A linear time algorithm for computing the Voronoi diagram of a convex polygon. In ACM [ACM87], pages 39–45. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p39-aggarwal/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p39-aggarwal/p39-aggarwal.

pdf. ACM order no. 508870.

Asano:1982:EDE

[AH82] Takao Asano and Tomio Hirata. Edge-deletion and edgecontraction problems. In ACM [ACM82], pages 245–254. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Adleman:1987:RPR

[AH87] L. Adleman and M. Huang. Recognizing primes in random polynomial time. In ACM [ACM87], pages 462–

469. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

                                        1987. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/28395/p462-adleman/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        28395/p462-adleman/p462-adleman.

pdf. ACM order no. 508870.

Adachi:1981:LLC

[AIK81] Akeo Adachi, Shigeki Iwata, and Takumi Kasai. Low level complexity for combinatorial games. In ACM [ACM81], pages 228–237. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Awerbuch:1984:FEC

[AIS84] B. Awerbuch, A. Israeli, and Y. Shiloach. Finding Euler circuits in logarithmic parallel time. In ACM [ACM84], pages 249–257. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Atallah:1982:GPM

[AK82] Mikhail J. Atallah and S. Rao Kosaraju. Graph problems on a mesh-connected processor array (preliminary version). In ACM [ACM82], pages 345–353. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Adleman:1988:USA

[AK88] Leonard Adleman and Kireeti Kompella. Using smoothness to achieve parallelism. In ACM [ACM88], pages 528–

538. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p528-adleman/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p528-adleman/p528-adleman.

pdf. ACM order no. 508880.

Ajtai:1983:SN

[AKS83] M. Ajtai, J. Komlós, and E. Szemerédi. An O(n log n) sorting network. In ACM [ACM83], pages 1–9. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Ajtai:1987:DSL

[AKS87] M. Ajtai, J. Komlos, and E. Szemeredi. Deterministic simulation in LOGSPACE. In ACM [ACM87], pages 132–

140. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p132-ajtai/;

http://www.acm.org/pubs/articles/
proceedings/stoc/28395/p132-
ajtai/p132-ajtai.pdf. ACM

order no. 508870.

Ajtai:1986:DSL

[AKSS86] M. Ajtai, J. Komlos, W. L. Steiger, and E. Szemeredi. Deterministic selection in O(log log N) parallel time. In ACM [ACM86], pages 188–195. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p188-ajtai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p188-
ajtai/p188-ajtai.pdf. ACM

order no. 508860.

Adleman:1986:FIP

[AL86] L. M. Adleman and H. W. Lenstra. Finding irreducible polynomials over finite fields. In ACM [ACM86], pages 350–

355. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p350-adleman/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p350-adleman/p350-adleman.

pdf. ACM order no. 508860.

Allender:1987:SCE

[All87]  E. Allender.  Some conse-

quences of the existence of pseudorandom generators. In ACM [ACM87], pages 151–

159. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/

stoc/28395/p151-allender/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p151-allender/p151-
allender.pdf. ACM order no.

508870.

Alon:1985:ESR

[Alo85] N. Alon. Expanders, sorting in rounds and superconcentrators of limited depth. In ACM [ACM85], pages 98–

102. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p98-alon/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/22145/p98-
alon/p98-alon.pdf. ACM or-

der no. 508850.

Alt:1984:CAF

[Alt84] Helmut Alt. Comparison of arithmetic functions with respect to boolean circuit depth. In ACM [ACM84], pages 466–

470. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Adler:1984:SAW

[AM84] Ilan Adler and Nimrod Megiddo.

A simplex algorithm whose average number of steps is bounded between two quadratic functions of the smaller dimension. In ACM [ACM84], pages 312–323. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Anderson:1985:PAM

[And85]  R. Anderson.  A paral-

lel algorithm for the maximal path problem. In ACM [ACM85], pages 33–37. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p33-anderson/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        22145/p33-anderson/p33-anderson.

pdf. ACM order no. 508850.

Angluin:1980:LGP

[Ang80]  Dana Angluin.  Local and

global properties in networks of processors (extended abstract). In ACM [ACM80], pages 82–

93. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Aho:1983:NIT

[AUY83] Alfred V. Aho, Jeffrey D. Ullman, and Mihalis Yannakakis. On notions of information transfer in VLSI circuits. In ACM [ACM83], pages 133–139. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Awerbuch:1984:ENS

[Awe84] Baruch Awerbuch. An efficient network synchronization protocol. In ACM [ACM84], pages 522–525. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Awerbuch:1987:ODA

[Awe87] B. Awerbuch. Optimal distributed algorithms for minimum weight spanning tree, counting, leader election, and related problems. In ACM [ACM87], pages 230–240. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. URL http:/

/www.acm.org/pubs/citations/
proceedings/stoc/28395/p230-
awerbuch/; http://www.acm.
org/pubs/articles/proceedings/
stoc/28395/p230-awerbuch/
p230-awerbuch.pdf. ACM or-

der no. 508870.

Awerbuch:1989:RDS

[Awe89] B. Awerbuch. Randomized distributed shortest paths algorithms. In ACM [ACM89], pages 490–500. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p490-awerbuch/
p490-awerbuch.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/73007/p490-

awerbuch/. ACM order no. 508890.

Babai:1985:TGT

[Bab85]  L. Babai.  Trading group

theory for randomness. In ACM [ACM85], pages 421–

429. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p421-babai/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p421-

                                        babai/p421-babai.pdf. ACM

order no. 508850.

Bach:1982:FAU

[Bac82] Eric Bach. Fast algorithms under the Extended Riemann Hypothesis: A concrete estimate. In ACM [ACM82], pages 290–295. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Bach:1983:HGR

[Bac83] Eric Bach. How to generate random integers with known factorization. In ACM [ACM83], pages 184–188. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Bach:1987:RAS

[Bac87] E. Bach. Realistic analysis of some randomized algorithms. In ACM [ACM87], pages 453–

461. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

                                        1987. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/28395/p453-bach/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/28395/p453-
                                        bach/p453-bach.pdf. ACM or-

der no. 508870.

Barrington:1986:BWP

[Bar86] D. A. Barrington. Boundedwidth polynomial-size branching programs recognize exactly those languages in NC^{1}. In ACM [ACM86], pages 1–5. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/12130/p1-barrington/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p1-barrington/p1-barrington.

pdf. ACM order no. 508860.

Blumer:1984:BCI

[BBE^{+}84] A. Blumer, J. Blumer, A. Ehrenfeucht, D. Haussler, and R. Mc-Connell. Building a complete inverted file for a set of text files in linear time. In ACM [ACM84], pages 349–358. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Baker:1983:AAM

[BBL83] Brenda S. Baker, Sandeep N.

Bhatt, and Frank Thomson Leighton. An approximation algorithm for Manhattan routing. In ACM [ACM83], pages 477–486. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Borodin:1980:TST

[BC80] A. Borodin and S. Cook. A time-space tradeoff for sorting on a general sequential model of computation. In ACM [ACM80], pages 294–

301. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Bhatt:1988:OSB

[BCHR88] Sandeep Bhatt, Fan Chung, Jia-Wei Hong, and Arnold Rosenberg. Optimal simulations by Butterfly Networks. In ACM [ACM88], pages 192–

204. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p192-bhatt/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p192-
bhatt/p192-bhatt.pdf. ACM

order no. 508880.

Ben-David:1989:TAC

[BDCG89] S. Ben-David, B. Chor, and O. Goldreich. On the theory of average case complexity. In ACM [ACM89], pages 204–216. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p204-ben-david/
p204-ben-david.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/73007/p204-

ben-david/. ACM order no. 508890.

Boyce:1982:FEP

[BDDG82] James E. Boyce, David P. Dobkin, Robert L. (Scot) Drysdale, and Leo J. Guibas. Finding extremal polygons. In ACM [ACM82], pages 282–289. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Borodin:1983:BWT

[BDFP83] Allan Borodin, Danny Dolev, Faith E. Fich, and Wolfgang Paul. Bounds for width two branching programs. In ACM [ACM83], pages 87–93. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Beame:1986:LPC

[Bea86] P. Beame. Limits on the power of concurrent-write parallel machines. In ACM [ACM86], pages 169–176. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p169-beame/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p169-
beame/p169-beame.pdf. ACM

order no. 508860.

Beame:1989:GST

[Bea89]  P. Beame.  A general se-

quential time-space tradeoff for finding unique elements. In ACM [ACM89], pages 197–

203. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p197-beame/p197-
beame.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p197-beame/. ACM

order no. 508890.

Blumer:1986:CLG

[BEHW86] A. Blumer, A. Ehrenfeucht, D. Haussler, and M. Warmuth. Classifying learnable geometric concepts with the Vapnik–Chervonenkis dimension. In ACM [ACM86], pages 273–282. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p273-blumer/;

                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p273-
                                        blumer/p273-blumer.pdf. ACM

order no. 508860.

Bern:1986:TPR

[Ber86] M. W. Bern. Two probabilistic results on rectilinear Steiner trees. In ACM [ACM86], pages 433–441. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p433-bern/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p433-
                                        bern/p433-bern.pdf. ACM or-

der no. 508860.

Broder:1984:EFT

[BFDS84] Andrei Broder, Michael Fischer, Danny Dolev, and Barbara Simons. Efficient fault tolerant routings in networks. In ACM [ACM84], pages 536–541. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Bollobas:1985:AFH

[BFF85] B. Bollobás, T. I. Fenner, and A. M. Frieze. An algorithm for finding Hamilton cycles in random graphs. In ACM [ACM85], pages 430–

439. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p430-bollobas/
                                        ; http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        22145/p430-bollobas/p430-

bollobas.pdf. ACM order no.

508850.

Baker:1985:SPT

[BFG85] B. S. Baker, S. Fortune, and E. Grosse. Stable prehension with three fingers. In ACM [ACM85], pages 114–

120. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p114-baker/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p114-
baker/p114-baker.pdf. ACM

order no. 508850.

Beeri:1981:PAD

[BFM^{+}81] Catriel Beeri, Ronald Fagin, David Maier, Alberto Mendelzon, Jeffrey Ullman, and Mihalis Yannakakis. Properties of acyclic database schemes. In ACM [ACM81], pages 355–

362. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Blum:1988:NIZ

[BFM88] Manuel Blum, Paul Feldman, and Silvio Micali. Noninteractive zero-knowledge and its applications. In ACM [ACM88], pages 103–112. ISBN 0-89791-264-0. LCCN QA 76.6

A13 1988. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/62212/p103-blum/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/62212/p103-

                                        blum/p103-blum.pdf. ACM or-

der no. 508880.

Babai:1982:IGB

[BGM82] László Babai, D. Yu. Grigoryev, and David M. Mount. Isomorphism of graphs with bounded eigenvalue multiplicity. In ACM [ACM82], pages 310–324. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Berkman:1989:HPP

[BGSV89] O. Berkman, Z. Galil, B. Schieber, and U. Vishkin. Highly parallelizable problems. In ACM [ACM89], pages 309–

319. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p309-berkman/p309-
                                        berkman.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/73007/p309-berkman/.

ACM order no. 508890.

Borodin:1982:RMS

[BH82] A. Borodin and J. E. Hopcroft. Routing, merging and sorting on parallel models of computation. In ACM [ACM82], pages 338–344. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Beame:1987:OBD

[BH87] P. Beame and J. Håstad. Optimal bounds for decision problems on the CRCW PRAM. In ACM [ACM87], pages 83–

93. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p83-beame/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/28395/p83-
beame/p83-beame.pdf. ACM

order no. 508870.

Birget:1989:PCR

[Bir89] J.-C. Birget. Proof of a conjecture of R. Kannan. In ACM [ACM89], pages 445–

453. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p445-birget/p445-
birget.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p445-birget/. ACM

order no. 508890.

Bent:1985:FMR

[BJ85] S. W. Bent and J. W.

John. Finding the median requires 2n comparisons. In ACM [ACM85], pages 213–

216. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p213-bent/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/22145/p213-
bent/p213-bent.pdf. ACM or-

der no. 508850.

Bentley:1984:SUE

[BJL^{+}84] J. L. Bentley, D. S. Johnson, F. T. Leighton, C. C. McGeoch, and L. A. McGeoch. Some unexpected expected behavior results for bin packing. In ACM [ACM84], pages 279–288. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Brent:1980:CCB

[BK80] R. P. Brent and H. T. Kung. The chip complexity of binary arithmetic. In ACM [ACM80], pages 190–200. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Blum:1989:DPC

[BK89] M. Blum and S. Kanna. Designing programs that check their work. In ACM [ACM89], pages 86–97. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p86-blum/p86-blum.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/73007/p86-blum/. ACM

order no. 508890.

Barringer:1984:NYM

[BKP84] Howard Barringer, Ruurd Kuiper, and Amir Pnueli. Now you may compose temporal logic specifications. In ACM [ACM84], pages 51–63. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Broder:1989:TST

[BKRU89] A. Z. Broder, A. R. Karlin, P. Raghavan, and E. Upfal. Trading space for time in undirected s–t connectivity. In ACM [ACM89], pages 543–549. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p543-broder/p543-
broder.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p543-broder/. ACM

order no. 508890.

Bhatt:1982:HAT

[BL82] Sandeep N. Bhatt and Charles E.

Leiserson. How to assemble tree machines (extended abstract). In ACM [ACM82], pages 77–

84. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Babai:1983:CLG

[BL83] László Babai and Eugene M.

Luks. Canonical labeling of graphs. In ACM [ACM83], pages 171–183. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Bruce:1985:PID

[BL85] K. B. Bruce and G. Longo.

Provable isomorphisms and domain equations in models of typed languages. In ACM [ACM85], pages 263–

272. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p263-bruce/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p263-
bruce/p263-bruce.pdf. ACM

order no. 508850.

Bloniarz:1980:SPA

[Blo80] Peter Bloniarz. A shortestpath algorithm with expected time O(n^{2} log n log^{∗} n). In ACM [ACM80], pages 378–

384. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Babai:1987:PGN

[BLS87a] L. Babai, E. Luks, and A. Seress. Permutation groups in NC. In ACM [ACM87], pages 409–

420. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

                                        1987. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/28395/p409-babai/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/28395/p409-
                                        babai/p409-babai.pdf. ACM

order no. 508870.

Borodin:1987:OOA

[BLS87b] A. Borodin, N. Linial, and M. Saks. An optimal online algorithm for metrical task systems. In ACM [ACM87], pages 373–382. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6

                                        A13 1987. URL http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/28395/p373-borodin/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        28395/p373-borodin/p373-borodin.

pdf. ACM order no. 508870.

Blum:1983:HES

[Blu83] Manuel Blum. How to exchange (secret) keys. In ACM [ACM83], pages 440–447. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Blum:1984:AME

[Blu84]  Norbert Blum.  An area-

maximum edge length tradeoff for VLSI layout. In ACM [ACM84], pages 92–97. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Blum:1989:AAC

[Blu89]  A. Blum.  An O(n^{0.4})-

approximation algorithm for 3coloring. In ACM [ACM89], pages 535–542. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p535-blum/p535-
blum.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p535-blum/. ACM

order no. 508890.

Bellare:1988:HSG

[BM88] Mihir Bellare and Silvio Micali.

How to sign given any trapdoor function. In ACM [ACM88], pages 32–42. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p32-bellare/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p32-
bellare/p32-bellare.pdf. ACM

order no. 508880.

Bertoni:1981:CCF

[BMS81] Alberto Bertoni, Giancarlo Mauri, and Nicoletta Sabadini.

A characterization of the class of functions computable in polynomial time on Random Access Machines. In ACM [ACM81], pages 168–176. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Bach:1984:SDP

[BMS84] Eric Bach, Gary Miller, and Jeffrey Shallit. Sums of divisors, perfect numbers, and factoring. In ACM [ACM84], pages 183–190. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Babai:1989:MPL

[BN89] L. Babai and N. Nisan. Multiparty protocols and logspacehard pseudorandom sequences. In ACM [ACM89], pages 1–11. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/73007/p1-
                                        babai/p1-babai.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        proceedings/stoc/73007/p1-

babai/. ACM order no. 508890.

Ben-Or:1983:LBA

[BO83] Michael Ben-Or. Lower bounds for algebraic computation trees. In ACM [ACM83], pages 80– 86. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Ben-Or:1988:DAS

[BO88] Michael Ben-Or. A deterministic algorithm for sparse multivariate polynomial interpolation. In ACM [ACM88], pages 301–309. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p301-ben-or/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p301-
ben-or/p301-ben-or.pdf. ACM

order no. 508880.

Ben-Or:1983:CSS

[BOCS83] Michael Ben-Or, Benny Chor, and Adi Shamir. On the cryptographic security of single RSA bits. In ACM [ACM83], pages 421–430. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Ben-Or:1986:FPA

[BOFKT86] M. Ben-Or, E. Feig, D. Kozen, and P. Tiwari. A fast parallel algorithm for determining all roots of a polynomial with real roots. In ACM [ACM86], pages 340–349. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p340-ben-or/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p340-
ben-or/p340-ben-or.pdf. ACM

order no. 508860.

Ben-Or:1988:MPI

[BOGKW88] Michael Ben-Or, Shafi Goldwasser, Joe Kilian, and Avi Widgerson. Multi-prover interactive proofs: how to remove intractability. In ACM [ACM88], pages 113–131. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p113-ben-or/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p113-
ben-or/p113-ben-or.pdf. ACM

order no. 508880.

Ben-Or:1984:CEA

[BOKR84] Michael Ben-Or, Dexter Kozen, and John Reif. The complexity of elementary algebra and geometry. In ACM [ACM84], pages 457–464. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Boppana:1984:TFB

[Bop84] Ravi B. Boppana. Threshold functions and bounded depth monotone circuits. In ACM [ACM84], pages 475–479. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Boppana:1989:OSB

[Bop89] R. B. Boppana. Optimal separations between concurrentwrite parallel machines. In ACM [ACM89], pages 320–

326. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p320-boppana/p320-
boppana.pdf;   http://www.
acm.org/pubs/citations/proceedings/

stoc/73007/p320-boppana/.

ACM order no. 508890.

Ben-Or:1988:CTN

[BOW88] Michael Ben-Or and Avi Wigderson. Completeness theorems for non-cryptographic faulttolerant distributed computation. In ACM [ACM88], pages 1–10. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p1-ben-or/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/62212/p1-
ben-or/p1-ben-or.pdf. ACM

order no. 508880.

Bilardi:1984:MAV

[BP84] G. Bilardi and F. P. Preparata.

A minimum area VLSI network for O(log n) time sorting. In ACM [ACM84], pages 64–70. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Bilardi:1987:STC

[BP87] G. Bilardi and F. P. Preparata.

Size-time complexity of Boolean networks for prefix computations. In ACM [ACM87], pages 436–442. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6

A13 1987. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/28395/p436-bilardi/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p436-bilardi/p436-bilardi.

pdf. ACM order no. 508870.

Bracha:1985:ERR

[Bra85] G. Bracha. An O(lg n) expected rounds randomized Byzantine generals protocol. In ACM [ACM85], pages 316–

326. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p316-bracha/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/22145/p316-
                                        bracha/p316-bracha.pdf. ACM

order no. 508850.

Brown:1980:KSA

[Bro80] Donna J. Brown. Kraft storage and access for list implementations(extended abstract). In ACM [ACM80], pages 100–

107. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Broder:1986:HHI

[Bro86] A. Z. Broder. How hard is it to marry at random? (O(n) the approximation of the permanent). In ACM [ACM86], pages 50–58. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p50-broder/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p50-
                                        broder/p50-broder.pdf. ACM

order no. 508860.

Borodin:1989:LBL

[BRT89] A. Borodin, W. L. Ruzzo, and M. Tompa. Lower bounds on the length of universal traversal sequences. In ACM [ACM89], pages 562–573. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p562-borodin/p562-
borodin.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p562-borodin/.

ACM order no. 508890.

Buss:1984:PPG

[BS84] Jonathan F. Buss and Peter W.

Shor. On the pagenumber of planar graphs. In ACM [ACM84], pages 98–100. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Bollobas:1985:EBD

[BS85] B. Bollobás and I. Simon. On the expected behavior of disjoint set union algorithms. In ACM [ACM85], pages 224–

231. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p224-bollobas/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p224-bollobas/p224-
bollobas.pdf. ACM order no.

508850.

Berman:1988:IFT

[BS88] Piotr Berman and J’anos Simon. Investigations of faulttolerant networks of computers. In ACM [ACM88], pages 66–77. ISBN 0-89791-264-0.

LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p66-berman/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/62212/p66-
                                        berman/p66-berman.pdf. ACM

order no. 508880.

Bshouty:1989:EDS

[Bsh89] N. H. Bshouty. On the extended direct sum conjecture. In ACM [ACM89], pages 177–

185. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p177-bshouty/p177-
                                        bshouty.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/73007/p177-bshouty/.

ACM order no. 508890.

Barrington:1987:FMF

[BT87] D. Barrington and D. Therien. Finite monoids and the fine structure of NC^{1}. In ACM [ACM87], pages 101–

109. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

                                        1987. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/28395/p101-barrington/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        28395/p101-barrington/p101-
                                        barrington.pdf. ACM order

no. 508870.

Buss:1985:PHF

[Bus85]  S. R. Buss.  The polyno-

mial hierarchy and fragments of bounded arithmetic. In ACM [ACM85], pages 285– 290. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p285-buss/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/22145/p285-
buss/p285-buss.pdf. ACM or-

der no. 508850.

Buss:1987:BFV

[Bus87] S. R. Buss. The Boolean formula value problem is in ALOG-TIME. In ACM [ACM87], pages 123–131. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6

A13 1987. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/28395/p123-buss/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/28395/p123-
buss/p123-buss.pdf. ACM or-

der no. 508870.

Bar-Yehuda:1982:AVC

[BYE82] R. Bar-Yehuda and S. Even. On approximating a vertex cover for planar graphs. In ACM [ACM82], pages 303–309. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Cai:1986:POR

[Cai86]  J. Y. Cai.  With proba-

bility one, a random oracle separates PSPACE from the polynomial-time hierarchy. In ACM [ACM86], pages 21–29. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p21-cai/;   http:

                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p21-
                                        cai/p21-cai.pdf. ACM order

no. 508860.

Canny:1988:SAG

[Can88] John Canny. Some algebraic and geometric computations in PSPACE. In ACM [ACM88], pages 460–469. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p460-canny/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/62212/p460-
                                        canny/p460-canny.pdf. ACM

order no. 508880.

Carter:1982:TST

[Car82] J. Lawrence Carter. The theory of signature testing for VLSI. In ACM [ACM82], pages 66–

76. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Chazelle:1980:DET

[CD80] Bernard Chazelle and David P. Dobkin. Detection is easier than computation (extended abstract). In ACM [ACM80], pages 146–153. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Cook:1982:BTP

[CD82] Stephen Cook and Cynthia Dwork. Bounds on the time for parallel RAM’s to compute simple functions. In ACM [ACM82], pages 231–233. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Chaum:1988:MUS

[CD88] David Chaum and Ivaa Damgard.

Multiparty unconditionally secure protocols. In ACM [ACM88], pages 11–19. ISBN 0-89791-264-0. LCCN QA 76.6

A13 1988. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/62212/p11-chaum/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/62212/p11-
chaum/p11-chaum.pdf. ACM

order no. 508880.

Coan:1985:DFS

[CDDS85] B. A. Coan, D. Dolev, C. Dwork, and L. Stockmeyer. The distributed firing squad problem. In ACM [ACM85], pages 335–

345. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p335-coan/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/22145/p335-
coan/p335-coan.pdf. ACM or-

der no. 508850.

Chazelle:1987:CCC

[CEG87] B. Chazelle, H. Edelsbrunner, and L. Guibas. The complexity of cutting convex polytypes. In ACM [ACM87], pages 66–

76. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p66-chazelle/;
 http://www.acm.org/pubs/

                                        articles/proceedings/stoc/
                                        28395/p66-chazelle/p66-chazelle.

pdf. ACM order no. 508870.

Chazelle:1989:LSC

[CEGS89] B. Chazelle, H. Edelsbrunner, L. Guibas, and M. Sharir. Lines in space-combinators, algorithms and applications. In ACM [ACM89], pages 382–

393. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p382-chazelle/
                                        p382-chazelle.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        proceedings/stoc/73007/p382-

chazelle/. ACM order no.

508890.

Chandra:1983:UFC

[CFL83a] Ashok K. Chandra, Steven Fortune, and Richard Lipton. Unbounded fan-in circuits and associative functions. In ACM [ACM83], pages 52–60. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Chandra:1983:MPP

[CFL83b] Ashok K. Chandra, Merrick L. Furst, and Richard J. Lipton. Multi-party protocols. In ACM [ACM83], pages 94–99. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Cosmadakis:1988:DOP

[CGKV88] Stavros Cosmadakis, Haim Gaifman, Paris Kanellakis, and Moshe Vardi. Decidable optimization problems for database logic programs. In ACM [ACM88], pages 477– 490. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p477-cosmadakis/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p477-cosmadakis/p477-
cosmadakis.pdf. ACM order

no. 508880.

Culik:1981:SEP

[CH81] Karel Culik and Tero Harju.

The ω-sequence equivalence problem for DOL systems is decidable. In ACM [ACM81], pages 1–6. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Chazelle:1981:CDP

[Cha81] Bernard M. Chazelle. Convex decompositions of polyhedra. In ACM [ACM81], pages 70–79. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Chaselle:1984:IET

[Cha84] Bernard Chaselle. Intersecting is easier than sorting. In ACM [ACM84], pages 125–134. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Chew:1981:UNF

[Che81] Paul Chew. Unique normal forms in term rewriting systems with repeated variables. In ACM [ACM81], pages 7–

18. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Chandra:1981:EBR

[CHMP81] Ashok Chandra, Joe Halpern, Albert Meyer, and Rohit Parikh. Equations between regular terms and an application to process logic. In ACM [ACM81], pages 384–390. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Chung:1985:SOS

[CHS85] F. R. Chung, D. J. Hajela, and P. D. Seymour. Selforganizing sequential search and Hilbert’s inequalities. In ACM [ACM85], pages 217–

223. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p217-chung/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/22145/p217-
                                        chung/p217-chung.pdf. ACM

order no. 508850.

Cosmadakis:1985:ETD

[CK85] S. S. Cosmadakis and P. C. Kanellakis. Equational theories and database constraints. In ACM [ACM85], pages 273–

284. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p273-cosmadakis/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p273-cosmadakis/p273-
cosmadakis.pdf. ACM order

no. 508850.

Chor:1989:ZOL

[CK89] B. Chor and E. Kushilevitz.

A zero-one law for Boolean privacy. In ACM [ACM89], pages 62–72. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p62-chor/p62-chor.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
stoc/73007/p62-chor/. ACM

order no. 508890.

Coffman:1986:PEA

[CL86] E. G. Coffman and F. T.

Leighton. A provably efficient algorithm for dynamic storage allocation. In ACM [ACM86], pages 77–90. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p77-coffman/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p77-
coffman/p77-coffman.pdf. ACM

order no. 508860.

Clarkson:1984:FET

[Cla84] Kenneth L. Clarkson. Fast expected-time and approximation algorithms for geometric minimum spanning trees. In ACM [ACM84], pages 342–348. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Clarkson:1985:PAP

[Cla85] K. Clarkson. A probabilistic algorithm for the post office problem. In ACM [ACM85], pages 175–184. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6

                                        A13 1985. URL http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/22145/p175-clarkson/
                                        ; http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        22145/p175-clarkson/p175-
                                        clarkson.pdf. ACM order no.

508850.

Clarkson:1986:FAR

[Cla86]  K. L. Clarkson.  Further

applications of random sampling to computational geometry. In ACM [ACM86], pages 414–423. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p414-clarkson/
                                        ; http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        12130/p414-clarkson/p414-
                                        clarkson.pdf. ACM order no.

508860.

Clarkson:1987:AAS

[Cla87]  K. Clarkson.  Approxima-

tion algorithms for shortest path motion planning. In ACM [ACM87], pages 56–65.

ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p56-clarkson/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p56-clarkson/p56-clarkson.

pdf. ACM order no. 508870.

Cleve:1986:LSC

[Cle86] R. Cleve. Limits on the security of coin flips when half the processors are faulty. In ACM [ACM86], pages 364–

369. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p364-cleve/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p364-
cleve/p364-cleve.pdf. ACM

order no. 508860.

Cleve:1988:CAF

[Cle88] Richard Cleve. Computing algebraic formulas with a constant number of registers. In ACM [ACM88], pages 254–

257. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p254-cleve/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p254-
cleve/p254-cleve.pdf. ACM

order no. 508880.

Chandra:1981:EID

[CLM81] Ashok K. Chandra, Harry R.

Lewis, and Johann A. Makowsky.

Embedded implicational dependencies and their inference problem. In ACM [ACM81], pages 342–354. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Chazelle:1981:MCV

[CM81] Bernard Chazelle and Louis Monier. A model of computation for VLSI with related complexity results. In ACM [ACM81], pages 318–

325. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Cunto:1984:ACS

[CM84] Walter Cunto and J. Ian Munro. Average case selection. In ACM [ACM84], pages 369–375. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Cohen:1989:SPT

[CM89] E. Cohen and N. Megiddo. Strongly polynomial-time and NC algorithms for detecting cycles in dynamic graphs. In ACM [ACM89], pages 523–

534. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p523-cohen/p523-
cohen.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p523-cohen/. ACM

order no. 508890.

Coppersmith:1984:ELG

[Cop84] Don Coppersmith. Evaluating logarithms in GF(2n). In ACM [ACM84], pages 201–207. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Chandra:1989:ERG

[CRRS89] A. K. Chandra, P. Raghavan, W. L. Ruzzo, and R. Smolensky. The electrical resistance of a graph captures its commute and cover times. In ACM [ACM89], pages 574–586. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p574-chandra/p574-
chandra.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p574-chandra/.

ACM order no. 508890.

Carlson:1980:GPM

[CS80] David A. Carlson and John E.

Savage. Graph pebbling with many free pebbles can be difficult. In ACM [ACM80], pages 326–332. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Carter:1985:CBS

[CSW85] L. Carter, L. Stockmeyer, and M. Wegman. The complexity of backtrack searches. In ACM [ACM85], pages 449–

457. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.

org/pubs/citations/proceedings/
stoc/22145/p449-carter/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p449-
carter/p449-carter.pdf. ACM

order no. 508850.

Cole:1984:HRP

[CSY84] Richard Cole, Micha Sharir, and Chee K. Yap. On k-hulls and related problems. In ACM [ACM84], pages 154–166. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Cook:1989:FIF

[CU89] S. Cook and A. Urquhart. Functional interpretations of feasibly constructive arithmetic. In ACM [ACM89], pages 107–

112. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p107-cook/p107-
cook.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p107-cook/. ACM

order no. 508890.

Culberson:1985:EUB

[Cul85] J. C. Culberson. The effect of updates in binary search trees. In ACM [ACM85], pages 205–

212. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p205-culberson/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p205-culberson/p205-

culberson.pdf. ACM order no.

508850.

Cole:1986:DCT

[CV86] R. Cole and U. Vishkin. Deterministic coin tossing and accelerating cascades: micro and macro techniques for designing parallel algorithms. In ACM [ACM86], pages 206– 219. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p206-cole/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/12130/p206-
cole/p206-cole.pdf. ACM or-

der no. 508860.

Courcoubetis:1986:RAF

[CVW86] C. Courcoubetis, M. Y. Vardi, and P. Wolper. Reasoning about fair concurrent programs. In ACM [ACM86], pages 283–294. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p283-courcoubetis/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p283-courcoubetis/p283-
courcoubetis.pdf. ACM order

no. 508860.

Coppersmith:1987:MMA

[CW87] D. Coppersmith and S. Winograd. Matrix multiplication via arithmetic progressions. In ACM [ACM87], pages 1–6. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        stoc/28395/p1-coppersmith/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        28395/p1-coppersmith/p1-coppersmith.

pdf. ACM order no. 508870.

Cypher:1980:AKP

[Cyp80] Allen Cypher. An approach to the k paths problem. In ACM [ACM80], pages 211–

217. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Bakker:1985:TSI

[dBMOZ85a] J. W. de Bakker, J. J. Meyer, E. R. Olderog, and J. I. Zucker. Transition systems, infinitary languages and the semantics of uniform concurrency. In ACM [ACM85], pages 252–

262. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p252-de_bakker/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        22145/p252-de_bakker/p252-
                                        de_bakker.pdf. ACM order no.

508850.

deBakker:1985:TSI

[dBMOZ85b] J. W. de Bakker, J. J. Meyer, E. R. Olderog, and J. I. Zucker. Transition systems, infinitary languages and the semantics of uniform concurrency. In ACM [ACM85], pages 252–

262. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13 1985. ACM order no. 508850.

deBakker:1982:DSC

[dBZ82] J. W. de Bakker and J. I.

Zucker. Denotational semantics of concurrency. In ACM [ACM82], pages 153–158. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Dolev:1983:SGG

[DDPW83] Danny Dolev, Cynthia Dwork, Nicholas Pippenger, and Avi Wigderson. Superconcentrators, generalizers and generalized connectors with limited depth. In ACM [ACM83], pages 42–51. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Dobkin:1986:PCP

[DEY86] D. Dobkin, H. Edelsbrunner, and C. K. Yap. Probing convex polytopes. In ACM [ACM86], pages 424–432. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p424-dobkin/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p424-
dobkin/p424-dobkin.pdf. ACM

order no. 508860.

Driscoll:1983:DPG

[DF83] James R. Driscoll and Merrick L. Furst. On the diameter of permutation groups. In ACM [ACM83], pages 152–160. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Dyer:1989:RPT

[DF89] M. Dyer and A. Frieze. A random polynomial time algorithm for approximating the volume of convex bodies. In ACM [ACM89], pages 375–

381. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p375-dyer/p375-
dyer.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p375-dyer/. ACM

order no. 508890.

deFraysseix:1988:SSS

[dFPP88] Hubert de Fraysseix, János Pach, and Richard Pollack. Small sets supporting fary embeddings of planar graphs. In ACM [ACM88], pages 426–

433. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/62212/p426-de_fraysseix/
p426-de_fraysseix.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/62212/p426-
de_fraysseix/. ACM order no.

508880.

Duris:1981:FTW

[DG81] Pavol Dūris̄ and Zvi Galil. Fooling a two-way automaton or one pushdown store is better than one counter for two way machines (preliminary version). In ACM [ACM81], pages 177–

188. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Duris:1982:TTB

[DG82] Pavol Dūris̄ and Zvi Galil.

Two tapes are better than one for nondeterministic machines. In ACM [ACM82], pages 1– 7. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Duris:1984:LBC

[DGS84] Pavol Duris, Zvi Galil, and Georg Schnitger. Lower bounds on communication complexity. In ACM [ACM84], pages 81–91. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Dolev:1988:TNA

[DGS88] Danny Dolev, Eli Gafni, and Nir Shavit. Toward a nonatomic era: l-exclusion as a test case. In ACM [ACM88], pages 78–92. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p78-dolev/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/62212/p78-
dolev/p78-dolev.pdf. ACM

order no. 508880.

Dolev:1984:PIA

[DHS84] Danny Dolev, Joe Halpern, and H. Raymond Strong. On the possibility and impossibility of achieving clock synchronization. In ACM [ACM84], pages 504–511. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Dolev:1984:NLF

[DHSS84] Danny Dolev, Joe Halpern, Barbara Simons, and Ray Strong. A new look at fault tolerant network routing. In ACM [ACM84], pages 526–535. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Dietz:1982:MOL

[Die82] Paul F. Dietz. Maintaining order in a linked list. In ACM [ACM82], pages 122–127. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Dolev:1981:OWB

[DKS^{+}81] Danny Dolev, Kevin Karplus, Alan Siegel, Alex Strong, and Jeffrey D. Ullman. Optimal wiring between rectangles. In ACM [ACM81], pages 312– 317. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

DeMillo:1980:CPE

[DL80] Richard A. DeMillo and Richard J. Lipton. The consistency of “P = NP”, and related problems with fragments of number theory. In ACM [ACM80], pages 45–57. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

DeMillo:1982:CP

[DLM82] Richard A. DeMillo, Nancy A.

Lynch, and Michael J. Merritt. Cryptographic protocols. In ACM [ACM82], pages 383–400. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Degano:1984:LPC

[DM84] Pierpaolo Degano and Ugo Montanari. Liveness properties as convergence in metric spaces. In ACM [ACM84], pages 31–38. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Dolev:1984:CFW

[DMMU84] Danny Dolev, David Maier, Ilarry Mairson, and Jeffrey Ullman. Correcting faults in write-once memory. In ACM [ACM84], pages 225–229. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Duris:1983:TNL

[DPGR83] Pavol Dūris̄, Wolfgang Paul, Zvi Galil, and Ruediger Reischuk. Two nonlinear lower bounds. In ACM [ACM83], pages 127–132. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Dwork:1986:FTN

[DPPU86] C. Dwork, D. Peleg, N. Pippenger, and E. Upfal. Fault tolerance in networks of bounded degree. In ACM [ACM86], pages 370–379. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p370-dwork/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p370-
dwork/p370-dwork.pdf. ACM

order no. 508860.

deRougemont:1984:UDF

[dR84] Michel de Rougemont. Uniform definability on finite structures with successor. In ACM [ACM84], pages 409–417. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Dolev:1982:PAM

[DS82] Danny Dolev and H. Raymond Strong. Polynomial algorithms for multiple processor agreement. In ACM [ACM82], pages 401–407. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Dietz:1987:TAM

[DS87] P. Dietz and D. Sleator. Two algorithms for maintaining order in a list. In ACM [ACM87], pages 365–

372. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p365-dietz/;
http://www.acm.org/pubs/articles/
proceedings/stoc/28395/p365-
dietz/p365-dietz.pdf. ACM

order no. 508870.

Dolev:1989:BCT

[DS89] D. Dolev and N. Shavit.

Bounded concurrrent timestamp systems are constructible. In ACM [ACM89], pages 454–

466. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p454-dolev/p454-
dolev.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p454-dolev/. ACM

order no. 508890.

Driscoll:1986:MDS

[DSST86] J. R. Driscoll, N. Sarnak, D. D.

Sleator, and R. E. Tarjan. Making data structures persistent. In ACM [ACM86], pages 109–121. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p109-driscoll/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p109-driscoll/p109-
driscoll.pdf. ACM order no.

508860.

Dymond:1983:SDM

[DT83] Patrick W. Dymond and Martin Tompa. Speedups of deterministic machines by synchronous parallel machines. In ACM [ACM83], pages 336–343. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Edelsbrunner:1986:TSA

[EG86] H. Edelsbrunner and L. J.

Guibas. Topologically sweeping an arrangement. In ACM [ACM86], pages 389– 403. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p389-edelsbrunner/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p389-edelsbrunner/p389-
edelsbrunner.pdf. ACM order

no. 508860.

Emerson:1982:DPE

[EH82] E. Allen Emerson and Joseph Y. Halpern. Decision procedures and expressiveness in the temporal logic of branching time. In ACM [ACM82], pages 169–180. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Ehrig:1980:CIL

[EM80] Hartmut Ehrig and Bernd Mahr. Complexity of implementations on the level of algebraic specifications. In ACM [ACM80], pages 281– 293. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Engelfriet:1983:IPA

[Eng83] Joost Engelfriet. Iterated pushdown automata and complexity classes. In ACM [ACM83], pages 365–373. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Emerson:1984:DBT

[ES84] E. Allen Emerson and A. Prasad Sistla. Deciding branching time logic. In ACM [ACM84], pages 14–24. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Fich:1985:OTT

[FadHRW85] F. E. Fich, F. Meyer auf der Heide, P. Ragde, and A. Wigderson. One, two, three … infinity: lower bounds for parallel computation. In ACM [ACM85], pages 48–58. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p48-fich/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/22145/p48-
fich/p48-fich.pdf. ACM or-

der no. 508850.

Fagin:1980:HCD

[Fag80] Ronald Fagin. Horn clauses and database dependencies (extended abstract). In ACM [ACM80], pages 123–

134. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Furedi:1986:CVD

[FB86] Z. Furedi and I. Barany. Computing the volume is difficult. In ACM [ACM86], pages 442–447. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/

                                        pubs/citations/proceedings/
                                        stoc/12130/p442-furedi/. ACM

order no. 508860.

Feder:1989:NFP

[Fed89]  T. Feder.  A new fixed

point approach for stable networks stable marriages. In ACM [ACM89], pages 513–

522. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p513-feder/p513-
                                        feder.pdf;   http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/73007/p513-feder/. ACM

order no. 508890.

Feidman:1983:DPP

[Fei83] Yishai A. Feidman. A decidable propositional probabilistic dynamic logic. In ACM [ACM83], pages 298–309. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Feldman:1985:FTM

[Fel85]  P. Feldman.  Fault toler-

ance of minimal path routings in a network. In ACM [ACM85], pages 327–

334. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p327-feldman/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        22145/p327-feldman/p327-feldman.

pdf. ACM order no. 508850.

Feldman:1986:NBN

[FFP86] P. Feldman, J. Friedman, and N. Pippenger. Non-blocking networks. In ACM [ACM86], pages 247–254. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p247-feldman/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p247-feldman/p247-feldman.

pdf. ACM order no. 508860.

Fiege:1987:ZKP

[FFS87] U. Fiege, A. Fiat, and A. Shamir.

Zero knowledge proofs of identity. In ACM [ACM87], pages 210–217. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6

A13 1987. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/28395/p210-fiege/;
http://www.acm.org/pubs/articles/
proceedings/stoc/28395/p210-
fiege/p210-fiege.pdf. ACM

order no. 508870.

Feder:1988:OAA

[FG88] Tomás Feder and Daniel Greene.

Optimal algorithms for approximate clustering. In ACM [ACM88], pages 434–

444. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p434-feder/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p434-
feder/p434-feder.pdf. ACM

order no. 508880.

Feldman:1982:PDL

[FH82] Yishai A. Feldman and David Harel. A probabilistic dynamic logic. In ACM [ACM82], pages 181–195. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Fich:1981:LBC

[Fic81] Faith E. Fich. Lower bounds for the cycle detection problem. In ACM [ACM81], pages 96– 105. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Fich:1983:NBP

[Fic83] Faith E. Fich. New bounds for parallel prefix circuits. In ACM [ACM83], pages 100–109. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Frederickson:1980:GSR

[FJ80] Greg N. Frederickson and Donald B. Johnson. Generalized selection and ranking (preliminary version). In ACM [ACM80], pages 420–428. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Franzblau:1984:ACR

[FK84] Deborah S. Franzblau and Daniel J. Kleitman. An algorithm for constructing regions with rectangles: Independence and minimum generating sets for collections of intervals. In ACM [ACM84], pages 167–174. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Friedman:1989:SER

[FKS89] J. Friedman, J. Kahn, and E. Szemerédi. On the second eigenvalue of random regular graphs. In ACM [ACM89], pages 587–598. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p587-friedman/
p587-friedman.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/73007/p587-

friedman/. ACM order no. 508890.

Frederickson:1984:ISC

[FL84] Greg N. Frederickson and Nancy A. Lynch. The impact of synchronous communication on the problem of electing a leader in a ring. In ACM [ACM84], pages 493–503. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Fellows:1989:SDE

[FL89] M. R. Fellows and M. A.

Langston. On search decision and the efficiency of polynomial-time algorithms. In ACM [ACM89], pages 501– 512. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p501-fellows/p501-
fellows.pdf;   http://www.

                                        acm.org/pubs/citations/proceedings/
                                        stoc/73007/p501-fellows/.

ACM order no. 508890.

Filotti:1980:PTA

[FM80] I. S. Filotti and Jack N. Mayer. A polynomial-time algorithm for determining the isomorphism of graphs of fixed genus. In ACM [ACM80], pages 236–

243. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Feldman:1988:OAB

[FM88] Paul Feldman and Silvio Micali. Optimal algorithms for Byzantine agreement. In ACM [ACM88], pages 148–

161. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p148-feldman/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        62212/p148-feldman/p148-feldman.

pdf. ACM order no. 508880.

Fiat:1989:IPS

[FN89] A. Fiat and M. Noar. Implicit O(1) probe search. In ACM [ACM89], pages 336–

344. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p336-fiat/p336-
                                        fiat.pdf;   http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/73007/p336-fiat/. ACM

order no. 508890.

Fiat:1988:SSM

[FNS^{+}88] Amos Fiat, Moni Naor, Alexandro Schäffer, Jeanette Schmidt, and Alan Siegel. Storing and searching a multikey table. In ACM [ACM88], pages 344–353. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p344-fiat/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/62212/p344-
fiat/p344-fiat.pdf. ACM or-

der no. 508880.

Fiat:1988:NOH

[FNSS88] Amos Fiat, Moni Naor, Jeanette Schmidt, and Alan Siegel. Non-oblivious hashing. In ACM [ACM88], pages 367–

376. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p367-fiat/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/62212/p367-
fiat/p367-fiat.pdf. ACM or-

der no. 508880.

Fortnow:1987:CPZ

[For87] L. Fortnow. The complexity of perfect zero-knowledge. In ACM [ACM87], pages 204–

209. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p204-fortnow/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/

                                        28395/p204-fortnow/p204-fortnow.

pdf. ACM order no. 508870.

Fischer:1980:OTL

[FP80] Michael J. Fischer and Michael S. Paterson. Optimal tree layout (preliminary version). In ACM [ACM80], pages 177–

189. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Friedl:1985:PTS

[FR85] K. Friedl and L. Rónyai. Polynomial time solutions of some problems of computational algebra. In ACM [ACM85], pages 153–162. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6

                                        A13 1985. URL http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/22145/p153-friedl/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/22145/p153-
                                        friedl/p153-friedl.pdf. ACM

order no. 508850.

Frederickson:1983:DSL

[Fre83] Greg N. Frederickson. Data structures for on-line updating of minimum spanning trees. In ACM [ACM83], pages 252–257. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Frederickson:1987:NAA

[Fre87] G. N. Frederickson. A new approach to all pairs shortest paths in planar graphs. In ACM [ACM87], pages 19–

28. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p19-frederickson/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p19-frederickson/p19-
frederickson.pdf. ACM order

no. 508870.

Frederickson:1984:DSL

[FS84] Greg N. Frederickson and Mandayam A. Srinivas. Data structures for on-line updating of matroid intersection solutions. In ACM [ACM84], pages 383–390. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Fortnow:1989:PCL

[FS89a] L. Fortnow and M. Sipser. Probabilistic computation and linear time. In ACM [ACM89], pages 148–156. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p148-fortnow/p148-
fortnow.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p148-fortnow/.

ACM order no. 508890.

Fredman:1989:CPC

[FS89b] M. Fredman and M. Saks. The cell probe complexity of dynamic data structures. In ACM [ACM89], pages 345–

354. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p345-fredman/p345-

                                        fredman.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/73007/p345-fredman/.

ACM order no. 508890.

Furer:1983:NFT

[FSS83] Martin Fürer, Walter Schnyder, and Ernst Specker. Normal forms for trivalent graphs and graphs of bounded valence. In ACM [ACM83], pages 161–170. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Fich:1985:PCE

[FT85] F. E. Fich and M. Tompa. The parallel complexity of exponentiating polynomials over finite fields. In ACM [ACM85], pages 38–47. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p38-fich/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/22145/p38-
                                        fich/p38-fich.pdf. ACM or-

der no. 508850.

Furer:1982:TDT

[Für82] Martin Fürer. The tight deterministic time hierarchy. In ACM [ACM82], pages 8–16. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Furer:1987:PRC

[Fur87] M. Furer. The power of randomness for communication complexity. In ACM [ACM87], pages 178–181. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. URL http://

www.acm.org/pubs/citations/
proceedings/stoc/28395/p178-
furer/;   http://www.acm.
org/pubs/articles/proceedings/
stoc/28395/p178-furer/p178-

furer.pdf. ACM order no. 508870.

Fagin:1985:ISM

[FV85] R. Fagin and M. Y. Vardi. An internal semantics for modal logic. In ACM [ACM85], pages 305–315. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6

A13 1985. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/22145/p305-fagin/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p305-
fagin/p305-fagin.pdf. ACM

order no. 508850.

Fortune:1988:PCM

[FW88] Steven Fortune and Gordon Wilfong. Planning constrained motion. In ACM [ACM88], pages 445–459. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p445-fortune/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p445-fortune/p445-fortune.

pdf. ACM order no. 508880.

Gabow:1983:ERT

[Gab83] Harold N. Gabow. An efficient reduction technique for degreeconstrained subgraph and bidirected network flow problems. In ACM [ACM83], pages 448–

456. ISBN 0-89791-099-0.

LCCN QA75.5.A14 1983. ACM order no. 508830.

Gacs:1983:RCC

[Gác83] Péter Gács. Reliable computation with cellular automata. In ACM [ACM83], pages 32– 41. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Galil:1984:OPA

[Gal84] Zvi Galil. Optimal parallel algorithms for string matching. In ACM [ACM84], pages 240–248. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Gabow:1984:SRT

[GBT84] Harold N. Gabow, Jon Louis Bentley, and Robert E. Tarjan. Scaling and related techniques for geometry problems. In ACM [ACM84], pages 135–143. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Gerth:1984:TLH

[Ger84] Rob Gerth. Transition logic: How to reason about temporal properties in a compositional way. In ACM [ACM84], pages 39–50. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Goodman:1985:SE

[GGMM85] J. Goodman, A. G. Greenberg, N. Madras, and P. March. On the stability of the Ethernet. In ACM [ACM85], pages 379– 387. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p379-goodman/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p379-goodman/p379-goodman.

pdf. ACM order no. 508850.

Gurevich:1982:TAG

[GH82] Yuri Gurevich and Leo Harrington. Trees, automata, and games. In ACM [ACM82], pages 60–65. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Goldwasser:1986:AAP

[GK86] S. Goldwasser and J. Kilian. Almost all primes can be quickly certified. In ACM [ACM86], pages 316–329. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p316-goldwasser/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p316-goldwasser/p316-
goldwasser.pdf. ACM order

no. 508860.

Galil:1986:NSP

[GKS86] Z. Galil, R. Kannan, and E. Szemeredi. On nontrivial separators for k-page graphs and simulations by nondeterministic one-tape Turing machines. In ACM [ACM86], pages 39–49. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p39-galil/;   http:
//www.acm.org/pubs/articles/

                                        proceedings/stoc/12130/p39-
                                        galil/p39-galil.pdf. ACM

order no. 508860.

Goldreich:1989:HCP

[GL89] O. Goldreich and L. A. Levin. A hard-core predicate for all one-way functions. In ACM [ACM89], pages 25–32. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p25-goldreich/
                                        p25-goldreich.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        proceedings/stoc/73007/p25-

goldreich/. ACM order no. 508890.

Gonnet:1981:LPS

[GM81] Gaston H. Gonnet and J. Ian Munro. A linear probing sort and its analysis (preliminary draft). In ACM [ACM81], pages 90–95. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Goldwasser:1982:PEH

[GM82] Shafi Goldwasser and Silvio Micali. Probabilistic encryption & how to play mental poker keeping secret all partial information. In ACM [ACM82], pages 365–377. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Goldwasser:1985:KCI

[GMR85] S. Goldwasser, S. Micali, and C. Rackoff. The knowledge complexity of interactive proof-systems. In ACM [ACM85], pages 291–

304. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p291-goldwasser/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p291-goldwasser/p291-
goldwasser.pdf. ACM order

no. 508850.

Goldberg:1984:FES

[GMS84] A. V. Goldberg and A. Marchetti-Spaccamela. On finding the exact solution of a zero-one knapsack problem. In ACM [ACM84], pages 359–368. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Goldreich:1987:HPM

[GMW87] O. Goldreich, S. Micali, and A. Wigderson. How to play ANY mental game. In ACM [ACM87], pages 218–

229. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p218-goldreich/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p218-goldreich/p218-
goldreich.pdf. ACM order no.

508870.

Goldwasser:1983:SSS

[GMY83] Shafi Goldwasser, Silvio Micali, and Andy Yao. Strong signature schemes. In ACM [ACM83], pages 431–439. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Goldreich:1987:TTS

[Gol87] O. Goldreich. Towards a theory of software protection and simulation by oblivious RAMs. In ACM [ACM87], pages 182–

194. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

                                        1987. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/28395/p182-goldreich/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        28395/p182-goldreich/p182-
                                        goldreich.pdf. ACM order no.

508870.

Gonnet:1984:DEE

[Gon84] Gaston H. Gonnet. Determining equivalence of expressions in random polynomial time. In ACM [ACM84], pages 334–341. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Galil:1981:EGP

[GP81] Zvi Galil and Wolfgang J. Paul. An efficient general purpose parallel computer. In ACM [ACM81], pages 247–

262. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Goldberg:1987:PSB

[GPS87] A. Goldberg, S. Plotkin, and G. Shannon. Parallel symmetrybreaking in sparse graphs. In ACM [ACM87], pages 315–

324. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p315-goldberg/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p315-goldberg/p315-
goldberg.pdf. ACM order no.

508870.

Goodman:1989:CRO

[GPS89] J. E. Goodman, R. Pollack, and B. Sturmfels. Coordinate representation of order types requires exponential storage. In ACM [ACM89], pages 405–410. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p405-goodman/p405-
goodman.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p405-goodman/.

ACM order no. 508890.

Gacs:1985:STD

[GR85]  P. Gács and J. Reif.  A

simple three-dimensional realtime cellular array. In ACM [ACM85], pages 388–

395. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p388-gacs/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/22145/p388-
gacs/p388-gacs.pdf. ACM or-

der no. 508850.

Greenberg:1982:TCB

[Gre82] Albert G. Greenberg. On the time complexity of broadcast communication schemes (preliminary version). In ACM [ACM82], pages 354–364. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Galil:1981:TSO

[GS81] Zvi Galil and Joel Seiferas. Time-space-optimal string matching (preliminary report). In ACM [ACM81], pages 106–

113. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Guibas:1983:PMG

[GS83] Leo J. Guibas and Jorge Stolfi. Primitives for the manipulation of general subdivisions and the computation of Voronoi diagrams. In ACM [ACM83], pages 221–234. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Goldberg:1985:CR

[GS85] A. Goldberg and M. Sipser. Compression and ranking. In ACM [ACM85], pages 440–

448. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p440-goldberg/
                                        ; http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        22145/p440-goldberg/p440-

goldberg.pdf. ACM order no.

508850.

Goldwasser:1986:PCV

[GS86] S. Goldwasser and M. Sipser.

Private coins versus public coins in interactive proof systems. In ACM [ACM86], pages 59–68. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p59-goldwasser/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p59-goldwasser/p59-
goldwasser.pdf. ACM order

no. 508860.

Gurevich:1988:NLT

[GS88] Yuri Gurevich and Saharon Shelah. Nondeterministic linear tasks may require substantially nonlinear deterministic time in the case of sublinear work space. In ACM [ACM88], pages 281–289. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p281-gurevich/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p281-gurevich/p281-
gurevich.pdf. ACM order no.

508880.

Gabow:1983:LTA

[GT83] Harold N. Gabow and Robert Endre Tarjan. A linear-time algorithm for a special case of disjoint set union. In ACM [ACM83], pages 246–251. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Goldberg:1986:NAM

[GT86] A. V. Goldberg and R. E. Tarjan. A new approach to the maximum flow problem. In ACM [ACM86], pages 136–146. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p136-goldberg/
                                        ; http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        12130/p136-goldberg/p136-
                                        goldberg.pdf. ACM order no.

508860.

Goldberg:1987:SMC

[GT87] A. Goldberg and R. Tarjan. Solving minimum-cost flow problems by successive approximation. In ACM [ACM87], pages 7–18. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6

                                        A13 1987. URL http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/28395/p7-goldberg/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/28395/p7-
                                        goldberg/p7-goldberg.pdf.

ACM order no. 508870.

Gabow:1988:AOS

[GT88a] Harold Gabow and Robert Tarjan. Almost-optimum speedups of algorithms for bipartite matching and related problems. In ACM [ACM88], pages 514–527. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/62212/p514-gabow/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p514-
gabow/p514-gabow.pdf. ACM

order no. 508880.

Goldberg:1988:FMC

[GT88b] Andrew Goldberg and Robert Tarjan. Finding minimum-cost circulations by canceling negative cycles. In ACM [ACM88], pages 388–397. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p388-goldberg/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p388-goldberg/p388-
goldberg.pdf. ACM order no.

508880.

Gabow:1988:FFG

[GW88] Harold Gabow and Herbert Westermann. Forests, frames, and games: algorithms for matroid sums and applications. In ACM [ACM88], pages 407–421. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p407-gabow/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p407-
gabow/p407-gabow.pdf. ACM

order no. 508880.

Guibas:1980:TSR

[GY80] Leo J. Guibas and F. Frances Yao. On translating a set of rectangles. In ACM [ACM80], pages 154–160. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Graham:1989:IRB

[GY89] R. L. Graham and A. C. Yao. On the improbability of reaching Byzantine agreements. In ACM [ACM89], pages 467–478. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p467-graham/p467-
                                        graham.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/73007/p467-graham/. ACM

order no. 508890.

Harel:1984:GRI

[Har84] David Harel. A general result on infinite trees and its applications. In ACM [ACM84], pages 418–427. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Harel:1985:LAF

[Har85] D. Harel. A linear algorithm for finding dominators in flow graphs and related problems. In ACM [ACM85], pages 185–

194. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p185-harel/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/22145/p185-
                                        harel/p185-harel.pdf. ACM

order no. 508850.

Haastad:1986:AOL

[Hås86] J. Håstad. Almost optimal lower bounds for small depth circuits. In ACM [ACM86], pages 6–20. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p6-hastad/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/12130/p6-
hastad/p6-hastad.pdf. ACM

order no. 508860.

Chan:1981:RCC

[hC81] Tat hung Chan. Reversal complexity of counter machines. In ACM [ACM81], pages 146–

157. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Hemachandra:1987:SEH

[Hem87] L. A. Hemachandra. The strong exponential hierarchy collapses. In ACM [ACM87], pages 110–

122. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p110-hemachandra/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p110-hemachandra/p110-
hemachandra.pdf. ACM order

no. 508870.

Hennessy:1984:MFP

[Hen84]  Matthew Hennessy.  Mod-

elling fair processes. In ACM [ACM84], pages 25–30. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Heath:1987:PNG

[HI87] L. Heath and S. Istrail. The page number of genus g graphs is O(g). In ACM [ACM87], pages 388–397. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. URL http://

www.acm.org/pubs/citations/
proceedings/stoc/28395/p388-
heath/;   http://www.acm.
org/pubs/articles/proceedings/
stoc/28395/p388-heath/p388-

heath.pdf. ACM order no. 508870.

Haastad:1989:FCU

[HL89] J. Håstad and T. Leighton. Fast computation using faulty hypercubes. In ACM [ACM89], pages 251–263. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p251-hastad/p251-
hastad.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p251-hastad/. ACM

order no. 508890.

Haastad:1987:RHP

[HLN87] J. Håstad, T. Leighton, and M. Newman. Reconfiguring a hypercube in the presence of faults. In ACM [ACM87], pages 274–284. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6

A13 1987. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/28395/p274-hastad/;
http://www.acm.org/pubs/articles/

proceedings/stoc/28395/p274-
hastad/p274-hastad.pdf. ACM

order no. 508870.

Haastad:1987:ABP

[HLR87] J. Håstad, T. Leighton, and B. Rogoff. Analysis of backoff protocols for multiple access channels. In ACM [ACM87], pages 241–

253. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p241-hastad/;
http://www.acm.org/pubs/articles/
proceedings/stoc/28395/p241-
hastad/p241-hastad.pdf. ACM

order no. 508870.

Halpern:1985:OPP

[HMM85] J. Y. Halpern, N. Megiddo, and A. A. Munshi. Optimal precision in the presence of uncertainty. In ACM [ACM85], pages 346–

355. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p346-halpern/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p346-halpern/p346-halpern.

pdf. ACM order no. 508850.

Hajnal:1988:CCG

[HMT88a] András Hajnal, Wolfgang Maass, and György Turán. On the communication complexity of graph properties. In ACM [ACM88], pages 186–191. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p186-hajnal/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/62212/p186-
                                        hajnal/p186-hajnal.pdf. ACM

order no. 508880.

Halpern:1988:KBA

[HMT88b] Joseph Halpern, Yjoram Moses, and Mark Tuttle. A knowledgebased analysis of zero knowledge. In ACM [ACM88], pages 132–147. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p132-halpern/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        62212/p132-halpern/p132-halpern.

pdf. ACM order no. 508880.

Hoffmann:1980:TIC

[Hof80] Christoph M. Hoffmann. Testing isomorphism of cone graphs(extended abstract). In ACM [ACM80], pages 244–251. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Halpern:1983:LRA

[HR83] Joseph Y. Halpern and Michael O. Rabin. A logic to reason about likelihood. In ACM [ACM83], pages 310–319. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Halstenberg:1988:DMC

[HR88] Bernd Halstenberg and Rudiger Reischuk. On different modes of communication. In ACM [ACM88], pages 162– 172. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p162-halstenberg/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p162-halstenberg/p162-
halstenberg.pdf. ACM order

no. 508880.

Heintz:1980:TPW

[HS80] J. Heintz and C. P. Schnorr.

Testing polynomials which are easy to compute (extended abstract). In ACM [ACM80], pages 262–272. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Hong:1982:NMN

[HS82] Zhu Hong and Robert Sedgewick.

Notes on merging networks (preliminary version). In ACM [ACM82], pages 296–302. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Hart:1984:PTL

[HS84a] Sergiu Hart and Micha Sharir.

Probabilistic temporal logics for finite and bounded models. In ACM [ACM84], pages 1–13. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Hochbaum:1984:PGP

[HS84b] Dorit S. Hochbaum and David B. Shmoys. Powers of graphs: A powerful approximation technique for bottleneck problems. In ACM [ACM84], pages 324–

333. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Haastad:1985:CST

[HS85] J. Håstad and A. Shamir. The cryptographic security of truncated linearly related variables. In ACM [ACM85], pages 356–

362. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p356-hastad/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/22145/p356-
                                        hastad/p356-hastad.pdf. ACM

order no. 508850.

Hartmanis:1983:SSN

[HSI83] J. Hartmanis, V. Sewelson, and N. Immerman. Sparse sets in NP-P: Exptime versus nexptime. In ACM [ACM83], pages 382–391. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Huang:1984:FPF

[Hua84] Ming-Deh A. Huang. Factorization of polynomials over finite fields and factorization of primes in algebraic number fields. In ACM [ACM84], pages 175–182. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Huang:1985:RHF

[Hua85]  M.-D. A. Huang.  Rie-

mann hypothesis and finding roots over finite fields. In ACM [ACM85], pages 121–

130. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p121-huang/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p121-
huang/p121-huang.pdf. ACM

order no. 508850.

Huynh:1985:CEP

[Huy85] D. T. Huynh. The complexity of the equivalence problem for commutative semigroups and symmetric vector addition systems. In ACM [ACM85], pages 405–

412. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p405-huynh/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p405-
huynh/p405-huynh.pdf. ACM

order no. 508850.

Halpern:1986:CRA

[HV86] J. Y. Halpern and M. Y.

Vardi. The complexity of reasoning about knowledge and time. In ACM [ACM86], pages 304–315. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p304-halpern/;
 http://www.acm.org/pubs/

                                        articles/proceedings/stoc/
                                        12130/p304-halpern/p304-halpern.

pdf. ACM order no. 508860.

Halpern:1988:RAK

[HV88] Joseph Halpern and Joseph Halpern Vardi. Reasoning about knowledge and time in asynchronous systems. In ACM [ACM88], pages 53–65. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p53-halpern/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/62212/p53-
                                        halpern/p53-halpern.pdf. ACM

order no. 508880.

Hawrusik:1981:CFC

[HVY81] F. Hawrusik, K. N. Venkataraman, and A. Yasuhara. Classes of functions for computing on binary trees (extended abstract). In ACM [ACM81], pages 19–27. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Ierardi:1989:QET

[Ier89] D. Ierardi. Quantifier elimination in the theory of an algebraically-closed field. In ACM [ACM89], pages 138–

147. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p138-ierardi/p138-
                                        ierardi.pdf;   http://www.

acm.org/pubs/citations/proceedings/
stoc/73007/p138-ierardi/.

ACM order no. 508890.

Ibarra:1980:CEP

[IL80] Oscar H. Ibarra and Brian S.

Leininger. The complexity of the equivalence problem for straight-line programs. In ACM [ACM80], pages 273–

280. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Impagliazzo:1989:PRG

[ILL89] R. Impagliazzo, L. A. Levin, and M. Luby. Pseudo-random generation from one-way functions. In ACM [ACM89], pages 12–24. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p12-impagliazzo/
p12-impagliazzo.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/73007/p12-
impagliazzo/. ACM order no.

508890.

Immerman:1982:RQC

[Imm82] Neil Immerman. Relational queries computable in polynomial time (extended abstract). In ACM [ACM82], pages 147–

152. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Immerman:1983:LWC

[Imm83] Neil Immerman. Languages which capture complexity classes.

In ACM [ACM83], pages 347–

354. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Impagliazzo:1989:LPC

[IR89] R. Impagliazzo and S. Rudich. Limits on the provable consequences of one-way permutations. In ACM [ACM89], pages 44–61. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p44-impagliazzo/
p44-impagliazzo.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/73007/p44-
impagliazzo/. ACM order no.

508890.

Iwano:1987:TCI

[IS87] K. Iwano and K. Steiglitz. Testing for cycles in infinite graphs with periodic structure. In ACM [ACM87], pages 46–

55. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p46-iwano/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/28395/p46-
iwano/p46-iwano.pdf. ACM

order no. 508870.

Istrail:1988:PUT

[Ist88] Sorin Istrail. Polynomial universal traversing sequences for cycles are constructible. In ACM [ACM88], pages 491–

503. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/

pubs/citations/proceedings/
stoc/62212/p491-istrail/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p491-istrail/p491-istrail.

pdf. ACM order no. 508880.

Inoue:1982:TDA

[ITT82] Katsushi Inoue, Itsuo Takanami, and Hiroshi Taniguchi. Twodimensional alternating Turing machines. In ACM [ACM82], pages 37–46. ISBN 0-89791-070-

2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Iwama:1983:UDS

[Iwa83] Kazuo Iwama. Unique decomposability of shuffled strings: A formal treatment of asynchronous time-multiplexed communication. In ACM [ACM83], pages 374–381. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

JaJa:1980:TST

[Ja’80]  Joseph Ja’Ja’.  Time-space

tradeoffs for some algebraic problems. In ACM [ACM80], pages 339–350. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Jimbo:1985:EOA

[JM85] S. Jimbo and A. Maruoka. Expanders obtained from affine transformations. In ACM [ACM85], pages 88–97. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13 1985. URL http:/

/www.acm.org/pubs/citations/

proceedings/stoc/22145/p88-
jimbo/;   http://www.acm.
org/pubs/articles/proceedings/
stoc/22145/p88-jimbo/p88-

jimbo.pdf. ACM order no. 508850.

Jerrum:1988:CRM

[JS88] Mark Jerrum and Alistair Sinclair. Conductance and the rapid mixing property for Markov chains: the approximation of permanent resolved. In ACM [ACM88], pages 235–

244. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p235-jerrum/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p235-
jerrum/p235-jerrum.pdf. ACM

order no. 508880.

Jia-wei:1980:SDS

[Jw80] Hong Jia-wei. On some deterministic space complexity problems. In ACM [ACM80], pages 310–317. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Jia-Wei:1981:CRB

[JWK81] Hong Jia-Wei and H. T. Kung. I/O complexity: The red-blue pebble game. In ACM [ACM81], pages 326–333. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Jia-Wei:1981:GAB

[JWR81] Hong Jia-Wei and Arnold L.

Rosenberg. Graphs that are almost binary trees (preliminary version). In ACM [ACM81], pages 334–341. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Joseph:1980:IRC

[JY80] Deborah Joseph and Paul Young. Independence results in Computer Science? (preliminary version). In ACM [ACM80], pages 58–69. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Joseph:1981:FPI

[JY81] Deborah Joseph and Paul Young. Fast programs for initial segments and polynomial time computation in weak models of arithmetic (preliminary abstract). In ACM [ACM81], pages 55–61. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Kim:1984:DDD

[KA84] Chul E. Kim and Timothy A.

Anderson. Digital disks and a digital compactness measure. In ACM [ACM84], pages 117–124.

ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Kosaraju:1986:OSB

[KA86] S. R. Kosaraju and M. J. Atallah. Optimal simulations between mesh-connected arrays of processors. In ACM [ACM86], pages 264–272. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p264-kosaraju/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p264-kosaraju/p264-
kosaraju.pdf. ACM order no.

508860.

Kaltofen:1982:PRM

[Kal82] Erich Kaltofen. A polynomial reduction from multivariate to bivariate integral polynomial factorization. In ACM [ACM82], pages 261–266. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Kaltofen:1985:CPG

[Kal85]  E. Kaltofen.  Computing

with polynomials given by straight-line programs I: greatest common divisors. In ACM [ACM85], pages 131–

142. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p131-kaltofen/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p131-kaltofen/p131-

kaltofen.pdf. ACM order no.

508850.

Kaltofen:1986:UCP

[Kal86]  E. Kaltofen.  Uniform clo-

sure properties of P-computable functions. In ACM [ACM86], pages 330–337. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p330-kaltofen/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p330-kaltofen/p330-
kaltofen.pdf. ACM order no.

508860.

Kaltofen:1987:SFH

[Kal87]  E. Kaltofen.  Single-factor

Hensel lifting and its application to the straight-line complexity of certain polynomials. In ACM [ACM87], pages 443–

452. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p443-kaltofen/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p443-kaltofen/p443-
kaltofen.pdf. ACM order no.

508870.

Kannan:1983:APN

[Kan83a] Ravi Kannan. Alternation and the power of nondeterminism. In ACM [ACM83], pages 344–346. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Kannan:1983:IAI

[Kan83b] Ravi Kannan. Improved algorithms for integer programming and related lattice problems. In ACM [ACM83], pages 193–206. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Karmarkar:1984:NPT

[Kar84]  N. Karmarkar.  A new

polynomial-time algorithm for linear programming. In ACM [ACM84], pages 302–311. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Kanellakis:1983:UID

[KCV83] Paris C. Kanellakis, Stavros S. Cosmadakis, and Moshe Y. Vardi. Unary inclusion dependencies have polynomial time inference problems. In ACM [ACM83], pages 264–277. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Kilian:1988:FCO

[Kil88] Joe Kilian. Founding cryptography on oblivious transfer. In ACM [ACM88], pages 20–31. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p20-kilian/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/62212/p20-
                                        kilian/p20-kilian.pdf. ACM

order no. 508880.

King:1981:MPA

[Kin81] K. N. King. Measures of parallelism in alternating computation trees (extended abstract). In ACM [ACM81], pages 189– 201. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

King:1988:LBC

[Kin88] Valerie King. Lower bounds on the complexity of graph properties. In ACM [ACM88], pages 468–476. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p468-king/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/62212/p468-
king/p468-king.pdf. ACM or-

der no. 508880.

Kissin:1982:MEC

[Kis82] Gloria Kissin. Measuring energy consumption in VLSI circuits: A foundation. In ACM [ACM82], pages 99–104. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Kannan:1980:OPD

[KL80a] Ravindran Kannan and Richard J.

Lipton. The orbit problem is decidable. In ACM [ACM80], pages 252–261. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Karp:1980:SCB

[KL80b] Richard M. Karp and Richard J. Lipton. Some connections between nonuniform and uniform complexity classes. In ACM [ACM80], pages 302–

309. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Kearns:1988:LPM

[KL88] Michael Kearns and Ming Li. Learning in the presence of malicious errors. In ACM [ACM88], pages 267–280. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p267-kearns/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/62212/p267-
                                        kearns/p267-kearns.pdf. ACM

order no. 508880.

Ko:1986:NOW

[KLD86] K. I. Ko, T. J. Long, and D. Z. Du. A note on one-way functions and polynomial-time isomorphisms. In ACM [ACM86], pages 295–303. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p295-ko/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p295-
                                        ko/p295-ko.pdf. ACM order

no. 508860.

Kannan:1984:PFN

[KLL84] R. Kannan, A. K. Lenstra, and L. Lovász. Polynomial factorization and nonrandomness of bits of algebraic and some transcendental numbers. In ACM [ACM84], pages 191–200. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Kleitman:1981:NLS

[KLLM81] Daniel Kleitman, Frank Thomson Leighton, Margaret Lepley, and Gary L. Miller. New layouts for the shuffle-exchange graph(extended abstract). In ACM [ACM81], pages 278–

292. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Koch:1989:WPE

[KLMR89] R. Koch, T. Leighton, B. Maggs, and S. Rao. Work-preserving emulations of fixed-connection networks. In ACM [ACM89], pages 227–240. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p227-koch/p227-
koch.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p227-koch/. ACM

order no. 508890.

Karp:1984:PAM

[KLMS84] Richard M. Karp, Michael Luby, and A. Marchetti-Spaccamela. A probabilistic analysis of multidimensional bin packing problems. In ACM [ACM84], pages 289–298. ISBN 0-89791-133-4.

LCCN QA 76.6 A13 1984. ACM

order no. 508840.

Kearns:1987:LBF

[KLPV87] M. Kearns, M. Li, L. Pitt, and L. Valiant. On the learnability of Boolean formulae. In ACM [ACM87], pages 285–

295. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

                                        1987. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/28395/p285-kearns/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/28395/p285-
                                        kearns/p285-kearns.pdf. ACM

order no. 508870.

Krishnamurthy:1981:EHT

[KM81] Balakrishnan Krishnamurthy and Robert N. Moll. Examples of hard tautologies in the propositional calculus. In ACM [ACM81], pages 28–37. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Kenyon-Mathieu:1989:VPO

[KMK89] C. Kenyon-Mathieu and V. King. Verifying partial orders. In ACM [ACM89], pages 367–

374. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p367-kenyon-mathieu/
                                        p367-kenyon-mathieu.pdf;
                                        http://www.acm.org/pubs/citations/
                                        proceedings/stoc/73007/p367-
                                        kenyon-mathieu/. ACM order

no. 508890.

Kurtz:1989:ICF

[KMR89] S. A. Kurtz, S. R. Mahaney, and J. S. Royer. The ismorphism conjecture fails relative to a random oracle. In ACM [ACM89], pages 157–166. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p157-kurtz/p157-
kurtz.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p157-kurtz/. ACM

order no. 508890.

Kannan:1988:IRG

[KNR88] Sampath Kannan, Moni Naor, and Steven Rudich. Implicit representation of graphs. In ACM [ACM88], pages 334–

343. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p334-kannan/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p334-
kannan/p334-kannan.pdf. ACM

order no. 508880.

Ko:1988:RPT

[Ko88] Ker-I. Ko. Relativized polynomial time hierarchies having exactly K levels. In ACM [ACM88], pages 245–

253. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p245-ko/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/62212/p245-

                                        ko/p245-ko.pdf. ACM order

no. 508880.

Kosaraju:1981:LSS

[Kos81] S. Rao Kosaraju. Localized search in sorted lists. In ACM [ACM81], pages 62–69. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Kosaraju:1982:DRV

[Kos82] S. Rao Kosaraju. Decidability of reachability in vector addition systems (preliminary version). In ACM [ACM82], pages 267–281. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Kosaraju:1986:PED

[Kos86] S. R. Kosaraju. Parallel evaluation of division-free arithmetic equations. In ACM [ACM86], pages 231–239. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p231-kosaraju/
                                        ; http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        12130/p231-kosaraju/p231-
                                        kosaraju.pdf. ACM order no.

508860.

Kozen:1983:PP

[Koz83] Dexter Kozen. A probabilistic PDL. In ACM [ACM83], pages 291–297. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Kozen:1984:PEE

[Koz84] Dexter Kozen. Pebblings, edgings, and equational logic. In ACM [ACM84], pages 428–435. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Klawe:1984:MFR

[KPPY84] Maria Klawe, Wolfgang J. Paul, Nicholas Pippenger, and Mihalis Yannakakis. On monotone formulae with restricted depth. In ACM [ACM84], pages 480–487. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Krizanc:1988:TRT

[KPU88] Danny Krizanc, David Peleg, and Eli Upfal. A timerandomness tradeoff for oblivious routing. In ACM [ACM88], pages 93–102. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p93-krizanc/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p93-
krizanc/p93-krizanc.pdf. ACM

order no. 508880.

Kim:1981:DSC

[KR81] Chul E. Kim and Azriel Rosenfeld. Digital straightness and convexity (extended abstract). In ACM [ACM81], pages 80–

89. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Karloff:1988:RAP

[KR88] Howard Karloff and Prabhakar Raghavan. Randomized algorithms and pseudorandom numbers. In ACM [ACM88], pages 310–321. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p310-karloff/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        62212/p310-karloff/p310-karloff.

pdf. ACM order no. 508880.

Krentel:1986:COP

[Kre86] M. W. Krentel. The complexity of optimization problems. In ACM [ACM86], pages 69–76. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p69-krentel/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p69-
                                        krentel/p69-krentel.pdf. ACM

order no. 508860.

Kahn:1984:EPG

[KS84] Jeff Kahn and Michael Saks. Every poset has a good comparison. In ACM [ACM84], pages 299–301. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Kalyanasundaram:1988:PWP

[KS88a] Balasubramanian Kalyanasundaram and George Schnitger. On the power of white pebbles. In ACM [ACM88], pages 258–266. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p258-kalyanasundaram/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p258-kalyanasundaram/
p258-kalyanasundaram.pdf.

ACM order no. 508880.

Kosaraju:1988:DCD

[KS88b] S. Rao Kosaraju and Gregory Sullivan. Detecting cycles in dynamic graphs in polynomial time. In ACM [ACM88], pages 398–406. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p398-kosaraju/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p398-kosaraju/p398-
kosaraju.pdf. ACM order no.

508880.

Kao:1989:LRG

[KS89] M.-Y. Kao and G. E. Shannon.

Local reorientation, global order, and planar topology. In ACM [ACM89], pages 286–

296. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p286-kao/p286-
kao.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p286-kao/. ACM

order no. 508890.

Karp:1980:LET

[KT80] Richard M. Karp and Robert Endre Tarjan. Linear expectedtime algorithms for connectivity problems (extended abstract). In ACM [ACM80], pages 368–

377. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Karlin:1986:PHE

[KU86] A. R. Karlin and E. Upfal. Parallel hashing — an efficient implementation of shared memory. In ACM [ACM86], pages 160–168. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p160-karlin/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p160-
                                        karlin/p160-karlin.pdf. ACM

order no. 508860.

Kurtz:1982:ROH

[Kur82] Stuart A. Kurtz. On the random oracle hypothesis. In ACM [ACM82], pages 224–230. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Karp:1985:SDP

[KUW85a] R. M. Karp, E. Upfal, and A. Wigderson. Are search and decision programs computationally equivalent? In ACM [ACM85], pages 464–

475. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p464-karp/;   http:

//www.acm.org/pubs/articles/
proceedings/stoc/22145/p464-
karp/p464-karp.pdf. ACM or-

der no. 508850.

Karp:1985:CPM

[KUW85b] R. M. Karp, E. Upfal, and A. Wigderson. Constructing a perfect matching is in random NC. In ACM [ACM85], pages 22–32. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p22-karp/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/22145/p22-
karp/p22-karp.pdf. ACM or-

der no. 508850.

Kapoor:1986:FAC

[KV86] S. Kapoor and P. M. Vaidya.

Fast algorithms for convex quadratic programming and multicommodity flows. In ACM [ACM86], pages 147–

159. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p147-kapoor/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p147-
kapoor/p147-kapoor.pdf. ACM

order no. 508860.

Kolaitis:1987:DPP

[KV87] P. Kolaitis and M. Vardi. The decision problem for the probabilities of higher-order properties. In ACM [ACM87], pages 425–435. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6

A13 1987. URL http://www.

                                        acm.org/pubs/citations/proceedings/
                                        stoc/28395/p425-kolaitis/
                                        ; http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        28395/p425-kolaitis/p425-
                                        kolaitis.pdf. ACM order no.

508870.

Kearns:1989:CLL

[KV89] M. Kearns and L. G. Valiant. Cryptographic limitations on learning Boolean formulae and finite automata. In ACM [ACM89], pages 433–

444. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p433-kearns/p433-
                                        kearns.pdf;   http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/73007/p433-kearns/. ACM

order no. 508890.

Karp:1984:FPA

[KW84] Richard M. Karp and Avi Wigderson. A fast parallel algorithm for the maximal independent set problem. In ACM [ACM84], pages 266–272. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Karchmer:1988:MCC

[KW88] Mauricio Karchmer and Avi Wigderson. Monotone circuits for connectivity require super-logarithmic depth. In ACM [ACM88], pages 539–

550. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/

stoc/62212/p539-karchmer/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p539-karchmer/p539-
karchmer.pdf. ACM order no.

508880.

Karp:1988:RPB

[KZ88] Richard Karp and Yanjun Zhang. A randomized parallel branch-and-bound procedure. In ACM [ACM88], pages 290–300. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p290-karp/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/62212/p290-
karp/p290-karp.pdf. ACM or-

der no. 508880.

Leivant:1981:CPP

[Lei81] Daniel Leivant. The complexity of parameter passing in polymorphic procedures. In ACM [ACM81], pages 38–45. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Leighton:1982:LSV

[Lei82] Frank Thomson Leighton. A layout strategy for VLSI which is provably good (extended abstract). In ACM [ACM82], pages 85–98. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Leighton:1984:TBC

[Lei84] Tom Leighton. Tight bounds on the complexity of parallel sorting. In ACM [ACM84], pages 71–80. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Lenstra:1983:FMP

[Len83] Arjen K. Lenstra. Factoring multivariate polynomials over finite fields. In ACM [ACM83], pages 189–192. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Levin:1984:PCA

[Lev84] Leonid A. Levin. Problems, complete in “average” instance. In ACM [ACM84], page 465. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Levin:1985:OWF

[Lev85] L. A. Levin. One-way functions and pseudorandom generators. In ACM [ACM85], pages 363–

365. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p363-levin/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/22145/p363-
                                        levin/p363-levin.pdf. ACM

order no. 508850.

Lichtenstein:1980:IGE

[Lic80] David Lichtenstein. Isomorphism for graphs embeddable on the projective plane. In ACM [ACM80], pages 218–

224. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Lloyd:1980:CPS

[Llo80] Errol L. Lloyd. Critical path scheduling of task systems with resource and processor constraints (extended abstract). In ACM [ACM80], pages 436–

446. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Lichtenstein:1987:IRS

[LLS87] D. Lichtenstein, N. Linial, and M. Saks. Imperfect random sources and discrete controlled processes. In ACM [ACM87], pages 169–

177. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p169-lichtenstein/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p169-lichtenstein/p169-
lichtenstein.pdf. ACM order

no. 508870.

Landau:1983:SRP

[LM83] Susan Landau and Gary Lee Miller. Solvability by radicals is in polynomial time. In ACM [ACM83], pages 140–151. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Leiserson:1985:ART

[LM85] C. E. Leiserson and F. M. Maley. Algorithms for routing and testing routability of planar VLSI layouts. In ACM [ACM85], pages 69–78. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

                                        1985. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/22145/p69-leiserson/
                                        ; http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        22145/p69-leiserson/p69-leiserson.

pdf. ACM order no. 508850.

Lueker:1988:MAD

[LM88] George Lueker and Mariko Molodowitch. More analysis of double hashing. In ACM [ACM88], pages 354–

359. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p354-lueker/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/62212/p354-
                                        lueker/p354-lueker.pdf. ACM

order no. 508880.

Lueker:1986:LPT

[LMR86] G. S. Lueker, N. Megiddo, and V. Ramachandran. Linear programming with two variables per inequality in poly-log time. In ACM [ACM86], pages 196–205. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p196-lueker/;
                                        http://www.acm.org/pubs/articles/

proceedings/stoc/12130/p196-
lueker/p196-lueker.pdf. ACM

order no. 508860.

Loebl:1988:LUS

[LN88] Martin Loebl and Jaroslav Nešetřil. Linearity and unprovability of set union problem. In ACM [ACM88], pages 360–366. ISBN 0-89791-264-

0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p360-loebl/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p360-
loebl/p360-loebl.pdf. ACM

order no. 508880.

Lubotzky:1986:EER

[LPS86] A. Lubotzky, R. Phillips, and P. Sarnak. Explicit expanders and the Ramanujan conjectures. In ACM [ACM86], pages 240–246. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p240-lubotzky/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p240-lubotzky/p240-
lubotzky.pdf. ACM order no.

508860.

Luby:1986:PRP

[LR86] M. Luby and C. Rackoff.

Pseudo-random permutation generators and cryptographic composition. In ACM [ACM86], pages 356–363. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/12130/p356-luby/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p356-
                                        luby/p356-luby.pdf. ACM or-

der no. 508860.

Lipton:1981:LBV

[LS81] Richard J. Lipton and Robert Sedgewick. Lower bounds for VLSI. In ACM [ACM81], pages 300–307. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Leighton:1986:TBM

[LS86] F. T. Leighton and P. Shor. Tight bounds for minimax grid matching, with applications to the average case analysis of algorithms. In ACM [ACM86], pages 91–103. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p91-leighton/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        12130/p91-leighton/p91-leighton.

pdf. ACM order no. 508860.

Lam:1989:TBC

[LTT89] T. Lam, P. Tiwari, and M. Tompa. Tradeoffs between communication and space. In ACM [ACM89], pages 217–

226. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p217-lam/p217-
                                        lam.pdf;   http://www.acm.

org/pubs/citations/proceedings/
stoc/73007/p217-lam/. ACM

order no. 508890.

Lubiw:1985:DLO

[Lub85a]  A. Lubiw.  Doubly lexical

orderings of matrices. In ACM [ACM85], pages 396–

404. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p396-lubiw/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p396-
lubiw/p396-lubiw.pdf. ACM

order no. 508850.

Luby:1985:SPA

[Lub85b] M. Luby. A simple parallel algorithm for the maximal independent set problem. In ACM [ACM85], pages 1–

10. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13 1985. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/22145/p1-
luby/p1-luby.pdf; http://
www.acm.org/pubs/citations/
proceedings/stoc/22145/p1-

luby/. ACM order no. 508850.

Landau:1986:IEP

[LV86] G. M. Landau and U. Vishkin.

Introducing efficient parallelism into approximate string matching and a new serial algorithm. In ACM [ACM86], pages 220–230. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p220-landau/;

                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p220-
                                        landau/p220-landau.pdf. ACM

order no. 508860.

Long:1983:HDD

[LW83] Douglas L. Long and Avi Wigderson. How discreet is the discrete log? In ACM [ACM83], pages 413–420. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Li:1986:NLB

[LY86] M. Li and Y. Yesha. New lower bounds for parallel computation. In ACM [ACM86], pages 177–187. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p177-li/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/12130/p177-
                                        li/p177-li.pdf. ACM order

no. 508860.

Lynch:1980:FAN

[Lyn80] Nancy A. Lynch. Fast allocation of nearby resources in a distributed system. In ACM [ACM80], pages 70–81. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Maass:1984:QLB

[Maa84] Wolfgang Maass. Quadratic lower bounds for deterministic and nondeterministic one-tape Turing machines. In ACM [ACM84], pages 401–408. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Manber:1984:MDI

[Man84] Udi Manber. On maintaining dynamic information in a concurrent environment. In ACM [ACM84], pages 273–278. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Mayr:1981:AGP

[May81] Ernst W. Mayr. An algorithm for the general Petri net reachability problem. In ACM [ACM81], pages 238–

246. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

MeyeraufderHeide:1985:FAD

[Mey85] F. Meyer auf der Heide. Fast algorithms for n-dimensional restrictions of hard problems. In ACM [ACM85], pages 413–

420. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p413-meyer_auf_

der_heide/. ACM order no. 508850.

Miller:1980:ITG

[Mil80] Gary Miller. Isomorphism testing for graphs of bounded genus. In ACM [ACM80], pages 225–

235. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Miller:1984:FSS

[Mil84] Gary L. Miller. Finding small simple cycle separators for 2connected planar graphs. In ACM [ACM84], pages 376–382. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Mirkowska:1980:CAA

[Mir80]  Grazyna Mirkowska.  Com-

plete axiomatization of algorithmic properties of program schemes with bounded nondeterministic interpretations. In ACM [ACM80], pages 14–21. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Mirzaian:1984:CRV

[Mir84] A. Mirzaian. Channel routing in VLSI. In ACM [ACM84], pages 101–107. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Manasse:1988:CAL

[MMS88] Mark Manasse, Lyle McGeoch, and Daniel Sleator. Competitive algorithms for on-line problems. In ACM [ACM88], pages 322–333. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p322-manasse/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/

62212/p322-manasse/p322-manasse.

pdf. ACM order no. 508880.

Motwani:1989:EGA

[Mot89] R. Motwani. Expanding graphs and the average-case analysis of algorithms for matchings and related problems. In ACM [ACM89], pages 550–

561. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p550-motwani/p550-
motwani.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p550-motwani/.

ACM order no. 508890.

Meyer:1980:DDL

[MP80] Albert R. Meyer and Rohit Parikh. Definability in dynamic logic. In ACM [ACM80], pages 1–7. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Miller:1987:NGT

[MR87] G. Miller and V. Ramachandran. A new graphy triconnectivity algorithm and its parallelization. In ACM [ACM87], pages 335–344. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. URL http://

www.acm.org/pubs/citations/
proceedings/stoc/28395/p335-
miller/;   http://www.acm.
org/pubs/articles/proceedings/
stoc/28395/p335-miller/p335-

miller.pdf. ACM order no. 508870.

Monien:1981:BCN

[MS81a] Burkhard Monien and Ivan Hal Sudborough. Bandwidth constrained NP-complete problems. In ACM [ACM81], pages 207– 217. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Muller:1981:PAG

[MS81b] David E. Muller and Paul E. Schupp. Pushdown automata, graphs, ends, second-order logic, and reachability problems. In ACM [ACM81], pages 46–54. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Mehlhorn:1982:VBT

[MS82] Kurt Mehlhorn and Erik M. Schmidt. Las Vegas is better than determinism in VLSI and distributed computing (extended abstract). In ACM [ACM82], pages 330–337. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Maass:1987:TTB

[MSS87] W. Maass, G. Schnitger, and E. Szemeredi. Two tapes are better than one for offline Turing machines. In ACM [ACM87], pages 94–100. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p94-maass/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/28395/p94-
maass/p94-maass.pdf. ACM

order no. 508870.

Manber:1982:PNA

[MT82] Udi Manber and Martin Tompa.

Probabilistic, nondeterministic, and alternating decision trees (preliminary version). In ACM [ACM82], pages 234–244. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Miller:1987:DPC

[MT87] G. Miller and S. Teng. Dynamic parallel complexity of computational circuits. In ACM [ACM87], pages 254–

263. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p254-miller/;
http://www.acm.org/pubs/articles/
proceedings/stoc/28395/p254-
miller/p254-miller.pdf. ACM

order no. 508870.

Mulmuley:1986:FPA

[Mul86] K. Mulmuley. A fast parallel algorithm to compute the rank of a matrix over an arbitrary field. In ACM [ACM86], pages 338–339. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p338-mulmuley/
; http://www.acm.org/pubs/

                                        articles/proceedings/stoc/
                                        12130/p338-mulmuley/p338-
                                        mulmuley.pdf. ACM order no.

508860.

Munro:1987:STK

[Mun87] J. I. Munro. Searching a two key table under a single key. In ACM [ACM87], pages 383–

387. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

                                        1987. URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/28395/p383-munro/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/28395/p383-
                                        munro/p383-munro.pdf. ACM

order no. 508870.

Mulmuley:1987:MEM

[MVV87] Ketan Mulmuley, Umesh V. Vazirani, and Vijay V. Vazirani. Matching is as easy as matrix inversion. In ACM [ACM87], pages 345–354. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. ACM order no. 508870.

Myers:1983:RAH

[Mye83] Dale Myers. The random access hierarchy. In ACM [ACM83], pages 355–364. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Natarajan:1987:LBF

[Nat87] B. K. Natarajan. On learning Boolean functions. In ACM [ACM87], pages 296–

304. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

                                        1987. URL http://www.acm.
                                        org/pubs/citations/proceedings/

stoc/28395/p296-natarajan/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p296-natarajan/p296-
natarajan.pdf. ACM order no.

508870.

Nisan:1989:CPD

[Nis89] N. Nisan. CREW PRAMS and decision trees. In ACM [ACM89], pages 327–335. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/73007/p327-
nisan/p327-nisan.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/73007/p327-

nisan/. ACM order no. 508890.

Naor:1987:FPA

[NNS87] J. Naor, M. Naor, and A. Schaffer. Fast parallel algorithms for chordal graphs. In ACM [ACM87], pages 355–

364. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p355-naor/;   http:
//www.acm.org/pubs/articles/
proceedings/stoc/28395/p355-
naor/p355-naor.pdf. ACM or-

der no. 508870.

Nivat:1982:ERM

[NP82] Maurice Nivat and Dominique Perrin. Ensembles reconnaissables de mots biinfinis. In ACM [ACM82], pages 47–59. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Naor:1989:UOW

[NY89] M. Naor and M. Yung. Universal one-way hash functions and their cryptographic applications. In ACM [ACM89], pages 33–43. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p33-naor/p33-naor.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/73007/p33-naor/. ACM

order no. 508890.

Orlitsky:1984:CSC

[OE84] Alon Orlitsky and Abbas El Gamal. Communication with secrecy constraints. In ACM [ACM84], pages 217–224. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Olderog:1983:CHL

[Old83] Ernst-Rüdiger Olderog. A characterization of Hoare’s logic for programs with Pascal-like procedures. In ACM [ACM83], pages 320–329. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Orlin:1981:CDL

[Orl81] James B. Orlin. The complexity of dynamic languages and dynamic optimization problems. In ACM [ACM81], pages 218– 227. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Orlin:1988:FSP

[Orl88] James Orlin. A faster strongly polynomial minimum cost flow algorithm. In ACM [ACM88], pages 377–387. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p377-orlin/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p377-
orlin/p377-orlin.pdf. ACM

order no. 508880.

Ong:1984:ESS

[OSS84] H. Ong, C. P. Schnorr, and A. Shamir. An efficient signature scheme based on quadratic equations. In ACM [ACM84], pages 208–216. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

oDunlaing:1983:RNA

[óSY83] Colm ó’Dúnlaing, Micha Sharir, and Chee K. Yap. Retraction: A new approach to motionplanning. In ACM [ACM83], pages 207–220. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Overmars:1980:DMC

[OvL80] Mark H. Overmars and Jan van Leeuwen. Dynamically maintaining configurations in the plane (detailed abstract). In ACM [ACM80], pages 135–

145. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Peleg:1985:CDL

[Pel85] D. Peleg. Concurrent dynamic logic. In ACM [ACM85], pages 232–239. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6

                                        A13 1985. URL http://www.
                                        acm.org/pubs/citations/proceedings/
                                        stoc/22145/p232-peleg/;
                                        http://www.acm.org/pubs/articles/
                                        proceedings/stoc/22145/p232-
                                        peleg/p232-peleg.pdf. ACM

order no. 508850.

Pippenger:1980:CSP

[Pip80] Nicholas Pippenger. Comparative schematology and pebbling with auxiliary pushdowns (preliminary version). In ACM [ACM80], pages 351–

356. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Pippenger:1982:PSP

[Pip82] Nicholas Pippenger. Probabilistic simulations (preliminary version). In ACM [ACM82], pages 17–26. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Pachl:1982:TPL

[PKR82] J. Pachl, E. Korach, and D. Rotem. A technique for proving lower bounds for distributed maximum-finding algorithms (preliminary version). In ACM [ACM82], pages 378–382. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Plaisted:1980:DIF

[Pla80] David A. Plaisted. On the distribution of independent formulae of number theory. In ACM [ACM80], pages 39–44. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Pnueli:1983:EFT

[Pnu83] Amir Pnueli. On the extremely fair treatment of probabilistic algorithms. In ACM [ACM83], pages 278–290. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Post:1984:MSE

[Pos84]  Mark J. Post.  Minimum

spanning ellipsoids. In ACM [ACM84], pages 108–116. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Parisi-Presicce:1981:FRE

[PP81] Francesco Parisi-Presicce. On the faithful regular extensions of iterative algebras. In ACM [ACM81], pages 368– 374. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Pan:1985:EPS

[PR85] V. Pan and J. Reif. Efficient parallel solution of linear systems. In ACM [ACM85], pages 143–152. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6

A13 1985. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/22145/p143-pan/. ACM

order no. 508850.

Pratt:1980:DAN

[Pra80] V. R. Pratt. Dynamic algebras and the nature of induction. In ACM [ACM80], pages 22–28. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Paterson:1981:BME

[PRS81] M. S. Paterson, W. L. Ruzzo, and L. Snyder. Bounds on minimax edge length for complete binary trees. In ACM [ACM81], pages 293–299. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Papadimitriou:1982:CC

[PS82] Christos H. Papadimitriou and Michael Sipser. Communication complexity. In ACM [ACM82], pages 196–200. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Paul:1980:ITA

[PSS80] Wolfgang J. Paul, Joel I. Seiferas, and Janos Simon. An information-theoretic approach to time bounds for on-line computation (preliminary version). In ACM [ACM80], pages 357–

367. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Pintz:1988:TIS

[PSS88] János Pintz, William Steiger, and Endre Szemerédi. Two infinite sets of primes with fast primality tests. In ACM [ACM88], pages 504–509. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p504-pintz/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p504-
pintz/p504-pintz.pdf. ACM

order no. 508880.

Peleg:1987:CDP

[PU87] D. Peleg and E. Upfal. Constructing disjoint paths on expander graphs. In ACM [ACM87], pages 264–273. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. URL http:/

/www.acm.org/pubs/citations/
proceedings/stoc/28395/p264-
peleg/;   http://www.acm.
org/pubs/articles/proceedings/
stoc/28395/p264-peleg/p264-

peleg.pdf. ACM order no. 508870.

Peleg:1988:TBS

[PU88] David Peleg and Eli Upfal. A tradeoff between space and efficiency for routing tables. In ACM [ACM88], pages 43–52. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/

                                        stoc/62212/p43-peleg/;   http:
                                        //www.acm.org/pubs/articles/
                                        proceedings/stoc/62212/p43-
                                        peleg/p43-peleg.pdf. ACM

order no. 508880.

Pitt:1989:MCD

[PW89] L. Pitt and M. K. Warmuth. The minimum consistent DFA problem cannot be approximated within and polynomial. In ACM [ACM89], pages 421–432. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p421-pitt/p421-
                                        pitt.pdf;   http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/73007/p421-pitt/. ACM

order no. 508890.

Papadimitriou:1982:CFS

[PY82] C. H. Papadimitriou and M. Yannakakis. The complexity of facets (and some facets of complexity). In ACM [ACM82], pages 255–260. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Papadimitriou:1988:OAC

[PY88a] Christos Papadimitriou and Mihalis Yannakakis. Optimization, approximation, and complexity classes. In ACM [ACM88], pages 229–234. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p229-papadimitriou/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        62212/p229-papadimitriou/

p229-papadimitriou.pdf. ACM

order no. 508880.

Papadimitriou:1988:TAI

[PY88b] Christos Papadimitriou and Mihalis Yannakakis. Towards an architecture-independent analysis of parallel algorithms. In ACM [ACM88], pages 510– 513. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p510-papadimitriou/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p510-papadimitriou/
p510-papadimitriou.pdf. ACM

order no. 508880.

Razborov:1989:MA

[Raz89]  A. A. Razborov.  On the

method of approximations. In ACM [ACM89], pages 167– 176. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p167-razborov/
p167-razborov.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/stoc/73007/p167-

razborov/. ACM order no. 508890.

Rabin:1989:VSS

[RBO89] T. Rabin and M. Ben-Or. Verifiable secret sharing and multiparty protocols with honest majority. In ACM [ACM89], pages 73–85. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/

                                        stoc/73007/p73-rabin/p73-
                                        rabin.pdf;   http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/73007/p73-rabin/. ACM

order no. 508890.

Reif:1980:LPP

[Rei80] John H. Reif. Logics for probabilistic programming (extended abstract). In ACM [ACM80], pages 8–13. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Reif:1982:SC

[Rei82]  John H. Reif.  Symmetric

complementation. In ACM [ACM82], pages 201–214. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Reif:1989:OSI

[Rei89] J. H. Reif. Optimal size integer division circuits. In ACM [ACM89], pages 264–

273. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

                                        URL http://www.acm.org/
                                        pubs/articles/proceedings/
                                        stoc/73007/p264-reif/p264-
                                        reif.pdf;   http://www.acm.
                                        org/pubs/citations/proceedings/
                                        stoc/73007/p264-reif/. ACM

order no. 508890.

Ronyai:1987:SAD

[Ron87] L. Ronyai. Simple algebras are difficult. In ACM [ACM87], pages 398–408. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. URL http://

www.acm.org/pubs/citations/
proceedings/stoc/28395/p398-
ronyai/;   http://www.acm.
org/pubs/articles/proceedings/
stoc/28395/p398-ronyai/p398-

ronyai.pdf. ACM order no. 508870.

Reif:1980:RM

[RS80] John H. Reif and Paul G.

Spirakis. Random matroids. In ACM [ACM80], pages 385– 397. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Reif:1981:DAS

[RS81] John Reif and Paul Spirakis.

Distributed algorithms for synchronizing interprocess communication within real time. In ACM [ACM81], pages 133– 145. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Rivest:1982:HRS

[RS82] Ronald L. Rivest and Adi Shamir. How to reuse a “writeonce” memory (preliminary version). In ACM [ACM82], pages 105–113. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Reif:1989:PNR

[RS89a] J. H. Reif and S. Sen. Polling: a new randomized sampling technique for computational geometry. In ACM [ACM89], pages 394–404. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p394-reif/p394-
reif.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p394-reif/. ACM

order no. 508890.

Rivest:1989:IFA

[RS89b] R. L. Rivest and R. E. Schapire. Inference of finite automata using homing sequences. In ACM [ACM89], pages 411–

420. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p411-rivest/p411-
rivest.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p411-rivest/. ACM

order no. 508890.

Ruzzo:1982:SBH

[RST82] Walter L. Ruzzo, Janos Simon, and Martin Tompa. Spacebounded hierarchies and probabilistic computations. In ACM [ACM82], pages 215–223. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Raghavan:1985:PGR

[RT85] P. Raghavan and C. D. Thompson. Provably good routing in graphs: regular arrays. In ACM [ACM85], pages 79–

87. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.

org/pubs/citations/proceedings/
stoc/22145/p79-raghavan/;
http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p79-raghavan/p79-raghavan.

pdf. ACM order no. 508850.

Reif:1983:LTS

[RV83] John H. Reif and Leslie G.

Valiant. A logarithmic time sort for linear size networks. In ACM [ACM83], pages 10–

16. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Sistla:1982:CPL

[SC82] A. P. Sistla and E. M. Clarke.

The complexity of propositional linear temporal logics. In ACM [ACM82], pages 159–168. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Seidel:1986:CHD

[Sei86] R. Seidel. Constructing higherdimensional convex hulls at logarithmic cost per face. In ACM [ACM86], pages 404–

413. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p404-seidel/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p404-
seidel/p404-seidel.pdf. ACM

order no. 508860.

Siegel:1986:AIF

[Sie86]  A. Siegel.  Aspects of in-

formation flow in VLSI circuits. In ACM [ACM86], pages 448–459. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p448-siegel/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p448-
siegel/p448-siegel.pdf. ACM

order no. 508860.

Simon:1981:SBP

[Sim81] Janos Simon. Space-bounded probabilistic Turing machine complexity classes are closed under complement (preliminary version). In ACM [ACM81], pages 158–167. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Sipser:1983:BSC

[Sip83a] Michael Sipser. Borel sets and circuit complexity. In ACM [ACM83], pages 61–69. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Sipser:1983:CTA

[Sip83b] Michael Sipser. A complexity theoretic approach to randomness. In ACM [ACM83], pages 330–335. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Smolensky:1987:AMT

[Smo87] R. Smolensky. Algebraic methods in the theory of lower bounds for Boolean circuit complexity. In ACM [ACM87], pages 77–82. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6

A13 1987. URL http://www.
acm.org/pubs/citations/proceedings/
stoc/28395/p77-smolensky/
; http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p77-smolensky/p77-smolensky.

pdf. ACM order no. 508870.

Suzuki:1985:MFP

[SNS85] H. Suzuki, T. Nishizeki, and N. Saito. Multicommodity flows in planar undirected graphs and shortest paths. In ACM [ACM85], pages 195–

204. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p195-suzuki/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p195-
suzuki/p195-suzuki.pdf. ACM

order no. 508850.

Spinrad:1983:TON

[Spi83] Jeremy Spinrad. Transitive orientation in 0(n2) time. In ACM [ACM83], pages 457–466. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Supowit:1980:HWP

[SPR80] Kenneth J. Supowit, David A.

Plaisted, and Edward M. Reingold. Heuristics for weighted perfect matching. In ACM [ACM80], pages 398–

419. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Sharir:1984:SPP

[SS84] Micha Sharir and Amir Schorr. On shortest paths in polyhedral spaces. In ACM [ACM84], pages 144–153. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Schnorr:1986:OSA

[SS86] C. P. Schnorr and A. Shamir. An optimal sorting algorithm for mesh connected computers. In ACM [ACM86], pages 255–263. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/12130/p255-schnorr/;
                                         http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        12130/p255-schnorr/p255-schnorr.

pdf. ACM order no. 508860.

Shelah:1987:TSR

[SS87] S. Shelah and J. Spencer. Threshold spectra for random graphs. In ACM [ACM87], pages 421–424. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. URL http://

                                        www.acm.org/pubs/citations/
                                        proceedings/stoc/28395/p421-
                                        shelah/;   http://www.acm.
                                        org/pubs/articles/proceedings/
                                        stoc/28395/p421-shelah/p421-

shelah.pdf. ACM order no.

508870.

Schmidt:1989:AUP

[SS89] J. P. Schmidt and A. Siegel. On aspects of university and performance for closed hashing. In ACM [ACM89], pages 355–366. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p355-schmidt/p355-
schmidt.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p355-schmidt/.

ACM order no. 508890.

Sleator:1981:DSD

[ST81] Daniel D. Sleator and Robert Endre Tarjan. A data structure for dynamic trees. In ACM [ACM81], pages 114–

122. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Sleator:1983:SAB

[ST83] Daniel Dominic Sleator and Robert Endre Tarjan. Selfadjusting binary trees. In ACM [ACM83], pages 235–245. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Sleator:1984:AEL

[ST84] Daniel Dominic Sleator and Robert Endre Tarjan. Amortized efficiency of list update rules. In ACM [ACM84], pages 488–492. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Storer:1980:NCM

[Sto80] James A. Storer. The node cost measure for embedding graphs on the planar grid (extended abstract). In ACM [ACM80], pages 201–210. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Stockmeyer:1983:CAC

[Sto83a] Larry Stockmeyer. The complexity of approximate counting. In ACM [ACM83], pages 118–126. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Stout:1983:TM

[Sto83b] Quentin F. Stout. Topological matching. In ACM [ACM83], pages 24–31. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Strong:1980:VEF

[Str80] H. R. Strong. Vector execution of flow graphs(extended abstract). In ACM [ACM80], pages 108–116. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Streett:1981:PDL

[Str81] Robert S. Streett. Propositional Dynamic Logic of looping and converse. In ACM [ACM81], pages 375–383. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Sleator:1986:RDT

[STT86] D. D. Sleator, R. E. Tarjan, and W. P. Thurston. Rotation distance, triangulations, and hyperbolic geometry. In ACM [ACM86], pages 122–135. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p122-sleator/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p122-sleator/p122-sleator.

pdf. ACM order no. 508860.

Sadri:1980:CAL

[SU80] Fereidoon Sadri and Jeffrey D.

Ullman. A complete axiomatization for a large class of dependencies in relational databases. In ACM [ACM80], pages 117–

122. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Safra:1989:ATL

[SV89] S. Safra and M. Y. Vardi.

On ω-automata and temporal logic. In ACM [ACM89], pages 127–137. ISBN 0-89791-307-

8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p127-safra/p127-
safra.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p127-safra/. ACM

order no. 508890.

Slot:1984:TVC

[SvEB84] C. Slot and P. van Emde Boas. On tape versus core an application of space efficient perfect hash functions to the invariance of space. In ACM [ACM84], pages 391–400. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Tompa:1980:OSW

[Tom80a] Martin Tompa. An optimal solution to a wire-routing problem (preliminary version). In ACM [ACM80], pages 161– 176. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Tompa:1980:TFT

[Tom80b] Martin Tompa. Two familiar transitive closure algorithms which admit no polynomial time, sublinear space implementations. In ACM [ACM80], pages 333–338. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Toueg:1980:DLF

[Tou80] Sam Toueg. Deadlock- and livelock-free packet switching networks. In ACM [ACM80], pages 94–99. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Turner:1983:PAB

[Tur83]  Jonathan Turner.  Proba-

bilistic analysis of bandwidth minimization algorithms. In ACM [ACM83], pages 467–476. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Tarjan:1986:LTA

[TV86] R. E. Tarjan and C. J. Van Wyk. A linear-time algorithm for triangulating simple polygons. In ACM [ACM86], pages 380–388. ISBN 0-89791-193-

8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p380-tarjan/;
http://www.acm.org/pubs/articles/
proceedings/stoc/12130/p380-
tarjan/p380-tarjan.pdf. ACM

order no. 508860.

Ukkonen:1980:DME

[Ukk80] Esko Ukkonen. A decision method for the equivalence of some non-real-time deterministic pushdown automata. In ACM [ACM80], pages 29–38. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Upfal:1984:PRB

[Upf84] Eli Upfal. A probabilistic relation between desirable and feasible, models of parallel computation. In ACM [ACM84], pages 258–265. ISBN 0-89791-133-4.

LCCN QA 76.6 A13 1984. ACM order no. 508840.

Upfal:1989:DPR

[Upf89] E. Upfal. An O(log N) deterministic packet routing scheme. In ACM [ACM89], pages 241–

249. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p241-upfal/p241-
upfal.pdf;   http://www.acm.
org/pubs/citations/proceedings/
stoc/73007/p241-upfal/. ACM

order no. 508890.

Ukkonen:1981:LKT

[USS81] Esko Ukkonen and Eljas Soisalon-Soininen. LALR(k) testing is PSPACE-complete. In ACM [ACM81], pages 202–

206. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Vaidya:1985:STT

[Vai85]  P. M. Vaidya.  Space-time

tradeoffs for orthogonal range queries. In ACM [ACM85], pages 169–174. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13 1985. URL http://

www.acm.org/pubs/citations/
proceedings/stoc/22145/p169-
vaidya/;   http://www.acm.
org/pubs/articles/proceedings/
stoc/22145/p169-vaidya/p169-

vaidya.pdf. ACM order no. 508850.

Vaidya:1987:ALP

[Vai87] P. M. Vaidya. An algorithm for linear programming which requires O(((m + n)n^{2} + (m + n)^{1.5}n)L) arithmetic operations. In ACM [ACM87], pages 29–

38. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p29-vaidya/;
http://www.acm.org/pubs/articles/
proceedings/stoc/28395/p29-
vaidya/p29-vaidya.pdf. ACM

order no. 508870.

Vaidya:1988:GHM

[Vai88] Pravin Vaidya. Geometry helps in matching. In ACM [ACM88], pages 422–425. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p422-vaidya/;
http://www.acm.org/pubs/articles/
proceedings/stoc/62212/p422-
vaidya/p422-vaidya.pdf. ACM

order no. 508880.

Valiant:1983:ELB

[Val83] L. G. Valiant. Exponential lower bounds for restricted monotone circuits. In ACM [ACM83], pages 110–117. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Valiant:1984:TL

[Val84] L. G. Valiant. A theory of the learnable. In ACM [ACM84], pages 436–445. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Vallee:1989:PFI

[Val89] B. Vallée. Provably fast integer factoring with quasi-uniform small quadratic residues. In ACM [ACM89], pages 98–

106. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989.

URL http://www.acm.org/
pubs/articles/proceedings/
stoc/73007/p98-vallee/p98-
vallee.pdf;   http://www.
acm.org/pubs/citations/proceedings/
stoc/73007/p98-vallee/. ACM

order no. 508890.

Vardi:1982:CRQ

[Var82] Moshe Y. Vardi. The complexity of relational query languages (extended abstract). In ACM [ACM82], pages 137–146. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Vazirani:1985:TSC

[Vaz85] U. V. Vazirani. Towards a strong communication complexity theory or generating quasirandom sequences from two communicating slightly-random sources. In ACM [ACM85], pages 366–378. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13 1985. URL http://

www.acm.org/pubs/citations/
proceedings/stoc/22145/p366-
vazirani/; http://www.acm.
org/pubs/articles/proceedings/
stoc/22145/p366-vazirani/
p366-vazirani.pdf. ACM or-

der no. 508850.

Vazirani:1987:ECU

[Vaz87] U. Vazirani. Efficiency considerations in using semi-random sources. In ACM [ACM87], pages 160–168. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. URL http://

www.acm.org/pubs/citations/
proceedings/stoc/28395/p160-
vazirani/; http://www.acm.
org/pubs/articles/proceedings/
stoc/28395/p160-vazirani/
p160-vazirani.pdf. ACM or-

der no. 508870.

Valiant:1981:USP

[VB81] L. G. Valiant and G. J. Brebner. Universal schemes for parallel communication. In ACM [ACM81], pages 263–

277. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Venkateswaran:1987:PCL

[Ven87] H. Venkateswaran. Properties that characterize LOGCFL. In ACM [ACM87], pages 141–

150. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13

1987. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/28395/p141-venkateswaran/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
28395/p141-venkateswaran/
p141-venkateswaran.pdf. ACM

order no. 508870.

Vishkin:1984:RSU

[Vis84]  Uzi Vishkin.  Randomized

speed-ups in parallel computation. In ACM [ACM84], pages 230–239. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Vitanyi:1982:RTS

[Vit82] Paul M. B. Vitányi. Real-time simulation of multicounters by oblivious one-tape Turing machines (preliminary draft). In ACM [ACM82], pages 27–36. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Vitanyi:1984:DEA

[Vit84] Paul M. B. Vitányi. Distributed elections in an archimedean ring of processors. In ACM [ACM84], pages 542–547. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Venkatesan:1988:RIG

[VL88] Ramarathnam Venkatesan and Leonid Levin. Random instances of a graph coloring problem are hard. In ACM [ACM88], pages 217–222. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        stoc/62212/p217-venkatesan/
                                        ;   http://www.acm.org/pubs/
                                        articles/proceedings/stoc/
                                        62212/p217-venkatesan/p217-
                                        venkatesan.pdf. ACM order

no. 508880.

Vardi:1985:IUL

[VS85] M. Y. Vardi and L. Stockmeyer. Improved upper and lower bounds for modal logics of programs. In ACM [ACM85], pages 240–

251. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p240-vardi/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p240-
vardi/p240-vardi.pdf. ACM

order no. 508850.

Valiant:1985:NED

[VV85a] L. G. Valiant and V. V. Vazirani. NP is as easy as detecting unique solutions. In ACM [ACM85], pages 458–

463. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p458-valiant/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p458-valiant/p458-valiant.

pdf. ACM order no. 508850.

Vazirani:1985:TPS

[VV85b] U. Vazirani and V. V. Vazirani. The two-processor scheduling problem is in R-NC. In ACM [ACM85], pages 11–21. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p11-vazirani/;
 http://www.acm.org/pubs/
articles/proceedings/stoc/
22145/p11-vazirani/p11-vazirani.

pdf. ACM order no. 508850.

Vardi:1984:ATT

[VW84] Moshe Y. Vardi and Pierre Wolper. Automata theoretic techniques for modal logics of programs: (extended abstract). In ACM [ACM84], pages 446–456. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

vonzurGathen:1983:PAA

[vzG83] Joachim von zur Gathen. Parallel algorithms for algebraic problems. In ACM [ACM83], pages 17–23. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Wigderson:1982:NAG

[Wig82] Avi Wigderson. A new approximate graph coloring algorithm. In ACM [ACM82], pages 325–329. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Willard:1982:MDS

[Wil82] Dan E. Willard. Maintaining dense sequential files in a dynamic environment (extended abstract). In ACM [ACM82], pages 114–121. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Willard:1984:LLP

[Wil84] Dan E. Willard. Log-logarithmic protocols for resolving Ethernet and semaphore conflicts. In ACM [ACM84], pages 512–521. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

Wilber:1985:WPH

[Wil85] R. Wilber. White pebbles help.

In ACM [ACM85], pages 103–

112. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13

1985. URL http://www.acm.
org/pubs/citations/proceedings/
stoc/22145/p103-wilber/;
http://www.acm.org/pubs/articles/
proceedings/stoc/22145/p103-
wilber/p103-wilber.pdf. ACM

order no. 508850.

Yannakakis:1981:ICD

[Yan81] Mihalis Yannakakis. Issues of correctness in database concurrency control by locking. In ACM [ACM81], pages 363–

367. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Yannakakis:1986:FPN

[Yan86] M. Yannakakis. Four pages are necessary and sufficient for planar graphs. In ACM [ACM86], pages 104–108. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/12130/p104-yannakakis/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
12130/p104-yannakakis/p104-
yannakakis.pdf. ACM order

no. 508860.

Yannakakis:1988:ECO

[Yan88] Mihalis Yannakakis. Expressing combinatorial optimization problems by linear programs. In ACM [ACM88], pages 223–228. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988.

URL http://www.acm.org/
pubs/citations/proceedings/
stoc/62212/p223-yannakakis/
;   http://www.acm.org/pubs/
articles/proceedings/stoc/
62212/p223-yannakakis/p223-
yannakakis.pdf. ACM order

no. 508880.

Yao:1980:EDP

[Yao80]  F. Frances Yao.  Efficient

dynamic programming using quadrangle inequalities. In ACM [ACM80], pages 429– 435. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Yao:1981:ELV

[Yao81a] Andrew C. Yao. The entropic limitations on VLSI computations(extended abstract). In ACM [ACM81], pages 308– 311. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Yao:1981:PCK

[Yao81b] Andrew Chi-Chih Yao. On the parallel computation for the knapsack problem. In ACM [ACM81], pages 123– 127. ISBN 0-89791-041-9 (paperback). LCCN ???? ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

Yao:1982:STT

[Yao82] Andrew C. Yao. Space-time tradeoff for answering range queries (extended abstract). In ACM [ACM82], pages 128–136. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

Yao:1983:SPA

[Yao83] F. Frances Yao. A 3-space partition and its applications. In ACM [ACM83], pages 258–263. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Yao:1989:CLC

[Yao89] A. C. Yao. Circuits and local computation. In ACM [ACM89], pages 186–196. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989. URL http://

www.acm.org/pubs/articles/
proceedings/stoc/73007/p186-
yao/p186-yao.pdf;   http:/
/www.acm.org/pubs/citations/
proceedings/stoc/73007/p186-

yao/. ACM order no. 508890.

Yap:1980:STT

[Yap80] C. K. Yap. Space-time tradeoffs and first order problems in a model of programs. In ACM [ACM80], pages 318– 325. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

Young:1983:SSP

[You83] Paul Young. Some structural properties of polynomial reducibilities and sets in NP. In ACM [ACM83], pages 392–401. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

Yao:1985:GAD

[YY85] A. C. Yao and F. F. Yao. A general approach to ddimensional geometric queries. In ACM [ACM85], pages 163– 168. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13 1985. URL http://

                                        www.acm.org/pubs/articles/
                                        proceedings/stoc/22145/p163-
                                        yao/p163-yao.pdf;   http:/
                                        /www.acm.org/pubs/citations/
                                        proceedings/stoc/22145/p163-

yao/. ACM order no. 508850.